### v13

Abstract
In an exchange economy composed of risk-averse individuals facing shocks, these shocks may become uninsurable if they are private information. Efficient shock smoothing across the economy requires self-reporting of the shocks that each individual privately suffers. This efficient allocation of resources leads to growing inequality between individuals, and the problem remains largely unsolved. This paper proposes a solution. The proposed mechanism (i) promises within-period full insurance by postponing risks ; (ii) does not lead to social ranking among the same age groups; and (iii) is sustainable under some conditions.
Key words: Incentives, optimal resource allocation, recursive contract, equality of same age group.
1 Introduction
In an exchange economy composed of risk-averse individuals facing shocks, these shocks may become uninsurable if they are private information. This problem has been studied extensively since the work of Green [1987]. These shocks are uninsurable because efficient shock smoothing across the economy requires self reporting of shocks that each individual privately suffers. Proposed solutions are limited. Inducing truth-telling at the expense of equality leads to immiserization, where the average expected lifetime utility of individuals converge to negative infinity while driving inequality without bound[Green, 1987, Thomas and Worrall, 1990]. If discount factor is approximately one, meaning individuals value the future and care little about the present, their contracts approach to the first best [Thomas and Worrall, 1990, Carrasco et al., 2019]. Atkeson and Lucas [1992] consider a scenario where discount factors represent resource prices, and social planners trade claims to current and future resources at these prices. They show that full information is required for efficient allocation.
This paper extends an alternative mechanism proposed by Marcet and Marimon [1992], which distributes risk across periods for the same individual. This approach could reduce the inequality observed in previous studies that allocate risk among individuals within each period. The mechanism proposed by Marcet and Marimon [1992] was an efficient growth mechanism combining state-contingent investment and transfers. This paper uses it for transfers in an exchange economy facing income shocks.
The first key finding of this paper is that our mechanism can achieve both efficiency and avoidance of immiserization under certain conditions   specifically, when individuals’ degree of risk aversion is not excessively high relative to the disparity between their highest and lowest possible incomes. Furthermore, the mechanism preserves the equality of opportunity within the same age group. Both efficiency and equality of opportunity are achieved by shifting risks to future periods. However, if the degree of risk aversion is too high relative to the disparity in income levels, challenges may arise. In such cases, it may become necessary to temporarily adjust the optimal transfer rules. This disruption occurs when the promised lifetime utility cannot be defined, given the assumption that each individual’s utility function is bounded from above. Nevertheless, even when the sufficient condition is not met, the mechanism continues to maintain the optimal allocation under the given condition and avoids immiserization.
The second key finding pertains to the potential sustainability of this mechanism in a society with a continuum of individuals, where the population, consisting of different age group, remains constant over time. In such a society, each individual faces a probability of dying in the next period, while new individuals continuously enter. Since risks exist across periods, maintaining unchanged expected lifetime utility across periods incurs costs. To ensure sustainability, earlier entrants must bear these costs, resulting in average consumption increasing with age. However, this does not lead to the formation of social ranking within the same age group. Importantly, this mechanism can ensure higher expected lifetime utility compared to autarky for all generations, while sustaining a positive balance of payments over time. Thus, a society that achieves both efficiency and equality of opportunity is feasible.
The paper is structured as follows. Section 2 introduces the notation and tarminology and proposes the mechanism for the inter-period risk transfers. Section 3 shows efficiency of the proposed mechanism and examines the properties for individuals. Section 4 investigates the potential sustainability of the mechanism. Section 5 provides some numerical examples. Section 6 evaluates the mechanism’s efficiency again using Bellman equation. Section 7 summarizes the conclusions.
2 Model
Consider an infinitely existing risk-neutral government for a society with a continuum of individuals in a unit interval. Each individual is risk-averse and will be alive in the next period with probability $\alpha\in(0,1)$. In each period, newly born risk-averse individuals enter the society, holding the total population constant. Individuals differ only in their periods of birth and death. Each individual has a discount factor $r\in(0,1)$. Let $\beta=\alpha r$ be the multiplication of $r$ and $\alpha$. A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game. Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $\{e_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ be a sequence of integrable random variables on $(\Omega,\mathcal{F},P)$ denoting identically and independently distributed income shocks of each individual in a sequence of periods. These income shocks take values in $E:=\{e^{1},e^{2},\dots,e^{M}\}$, where $e^{i}\in\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$. The government considers a transfer mechanism $\mathcal{M}$ that repeats from period zero to infinity, using a sequence of transfers (controls) $\{\tau_{t}\}_{t=0}^{\infty}$, $\tau_{t}\in\mathbb{R}$ and a sequence of utility control variables $\{\lambda_{t}\}_{t=0}^{\infty}$, $\lambda_{t}\in\mathbb{R}_{++}$ which will be specified later.
2.1 Pareto optimal transfer mechanism with a utility control variable $\lambda_{0}$
First, we derive an optimal transfer function which, under symmetric income information, ensures a level of utility corresponding to a utility control variable $\lambda_{0}>0$ Following Marcet and Marimon [1992], we will use this utility control variable $\lambda_{0}$ as a state variable under asymmetric income information. For this purpose, in this subsection only, we consider the case where individuals cannot hide their income from the government in each period.
In period 00, the government is considering an efficient transfer mechanism that increases individuals’ lifetime utility as much as possible, but the government cannot allow an infinite net subsidy. The government solves the following problem, having decided on an upper bound for the net subsidy.
()$(\ast)\begin{cases}\quad\max_{\{\tau_{t}\}_{t=0}^{\infty}}(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})\right]\\ \text{subject to}\\ \quad(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}(\tau_{t})\right]\leq\text{Const},\end{cases}$
where $\mathrm{E}_{0}$ is the conditional expectation given $e_{0}$, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period. The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
Without specifying the upper bound of the net subsidy, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines the upper bound of the net subsidy as well. Thus, the problem ()($\ast$)( ∗ ) can be rewritten as the following problem, where the objective function is the Lagrangian of problem ()($\ast$)( ∗ ) with an undetermined multiplier $01/\lambda_{0}>$ 0.
$\max_{\{\tau_{t}\}_{t=0}^{\infty}}$ $(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}[\lambda_{0}u(c_{t})-\tau_{t}]\right].$ (1)
From the first order condition,
$u^{\prime}(c_{t})=\frac{1}{\lambda_{0}}\text{ for }t\in\mathbb{N}\cup\{0\}$
and we obtain $c_{t}=(u^{\prime})^{-1}(\lambda_{0}^{-1})$ and $\tau_{t}=(u^{\prime})^{-1}(\lambda_{0}^{-1})-e_{t}$ for $t\in\mathbb{N}\cup\{0\}$. This is the Pareto optimal transfer mechanism that gives the individual constant consumption in each period.
Let $v_{1}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing each individual’s lifetime utility provided by the transfer mechanism. Let $v_{2}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing the expected balance of transfers. Then, for given $\lambda_{0}>0$ and $e_{0}\in E$, these values are written as follows.
$v_{1}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathrm{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})]=u((u^{\prime})^{-1}(\lambda_{0}^{-1}))=:\overline{v}_{1}(\lambda_{0})$
$v_{2}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathrm{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}(-\tau_{t})].$
Note that the function $v_{1}$ could be seen as a function of the value $\lambda_{0}$ whose inverse was the Lagrangian multiplier. We write it as $\overline{v}_{1}$. The function $\overline{v}_{1}$ is bijective and increasing. On the other hand, we see that $v_{2}(\lambda_{0},e_{0})$ depends on income. So we write its expectation as $\overline{v}_{2}(\lambda_{0})$.
$\overline{v}_{2}(\lambda_{0})$ $:=\mathrm{E}[v_{2}(\lambda_{0},e_{t})].$
In Section 4 the government will adjust the value of $\lambda_{0}$ so that the expected balance of total transfers across generations is non-negative.
Let $\tau^{*}(\cdot\ |\lambda_{0})$ be a time-invariant transfer function, which is the solution to the problem (1). That is, $\tau^{*}:\mathbb{R}_{+}\times\mathbb{R}_{++}\to\mathbb{R}_{+}$ is defined such that, for a reported income $x\in\mathbb{R}_{+}$ and a utility control variable $\lambda>0$
$\tau^{*}(x|\lambda):=(u^{\prime})^{-1}(\lambda^{-1})-x.$
We will use these notation in the following section for the case where income is private information. There we observe that a utility control variable $\lambda_{t}$ can play the role of a state variable and the role of an incentive device. Let $\mathcal{T}$ be the set of all admissible control functions there, and let $\tau:\mathbb{R}_{+}\times\mathbb{R}_{++}\to\mathbb{R}_{+}$ be a typical element of $\mathcal{T}$. Then the lifetime utility of individuals who are guaranteed utility level $\lambda_{t}$ at an initial period $t$ is given by
$J(\lambda_{t},t,\tau)=(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T},$ (2)
where $\mathrm{E}_{t}$ denotes the conditional expectation given $e_{t}$. We will see that $\tau^{*}\in\mathcal{T}$ could be an optimal control function for any initial period $t$ and for later specified $\{\lambda_{k}\}_{k=t}^{\infty}$ such that
$J(\lambda_{t},t,\tau^{*})=\sup_{\tau\in\mathcal{T}}$ $(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right].$ (3)
2.2 Pareto optimal transfer mechanism with asymmetric information
If income is private information, the government has to ask individuals to declare their income. In general, a mechanism $\Gamma$ that arranges transfers $\{\tau_{t}\}_{t=0}^{\infty}$ needs to satisfy sequential incentive compatibility defined below. For this definition, let $z$ denote the current observable state, which can be a finite vector of state variables, and may include the past history of observables or be in a recursive form. The distribution of the tomorrow’s state $z^{\prime}$ is determined by this $z$ and the current level of consumption $x$. We write $\pi(x)$ for the value of each individual’s current payoff and $v(z)$ for the present value of the mechanism $\Gamma$ at $z$.
Definition 1.
[Sequential incentive compatibility, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be sequentially incentive compatible if for every sate $z$
$v(z)=(1-\beta)\pi(x)+\beta\mathrm{E}_{(x,z)}[v(z^{\prime})]\geq(1-\beta)\pi(\tilde{x})+\beta\mathrm{E}_{(\tilde{x},z)}[v(z^{\prime})],$
where $E_{(x,z)}$ means that the conditional expectation given $(x,z)$.
We further evaluate a mechanism according to the following notion.
Definition 2.
[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequnetially incentive compatible mechanism.
In addition to taking these things into account, since it is known to lead to an immisarization outcome, the government does not hope to manipulate the transfer function $\tau^{*}$. The government takes an alternative approach. It is to vary the utility control variable $\lambda$ in each period. This could be an incentive to tell the truth. For a given $\lambda$ in a given period, the government promises to ensure that the lifetime utility level $\overline{v}_{1}(\lambda)$, and the utility control variable for the next period is renewed taking into account the current utility control variable and the income reported in the current period. If an individual truthfully reports income $e$ in a period $t$, then the lifetime utility from that period should be
$\overline{v}_{1}(\lambda)=\mathrm{E}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]},$ (4)
where $\lambda^{\prime}(\cdot|\lambda):E\to\mathbb{R}_{++}$ is the utility control variable for the next period. If $\lambda^{\prime}$ satisfies (4), then $\mathrm{E}[\overline{v}_{1}(\lambda^{\prime}(e|\lambda)]=\overline{v}_{1}(\lambda)$. Truth telling is then induced if, given the current utility control variable $\lambda$, the renewed utility control variable $\lambda^{\prime}$ satisfies the following incentive constraints.
$(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))$
$\geq(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\text{ for }e,\tilde{e}\in E.$ (5)
Such an utility control variable $\lambda^{\prime}$ that satisfies both (4) and (5) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$. By a feasible $\lambda^{\prime}$ we mean that the updated utility control variable keeps the level of expected utility for the next period below the supremum of each individual’s utility function. That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$. For given $\lambda>0$, if $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))<S$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))\big{)}.$ (6)
Since $v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)=(1-\beta)(e-\mathrm{E}[e_{t}])$, if the reported income $e\in E$ is above the average income, $\mathrm{E}[e_{t}]$, this will be rewarded with higher utility control variables for the next period. Conversely, if the reported income $e\in E$ is below the average income, $\mathrm{E}[e_{t}]$, this will result in lower utility control variables for the next period. Note that $u(e+\tau^{*}(e|\lambda))=u((u^{\prime})^{-1}(\lambda^{-1}))=\overline{v}_{1}(\lambda)$. From this we see that $\lambda^{\prime}$ defined in (6) satisfies (4):
$\mathrm{E}[(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))]$
$=\mathrm{E}[\overline{v}_{1}(\lambda)+\lambda^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))]=\overline{v}_{1}(\lambda).$
Furthermore, remind that the government’s objective function in problem (1) with $\lambda_{0}=\lambda$:
$\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)$
is optimal at true income report $e$. It follows that
$\lambda\big{[}(1-\beta)$ $u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)$
$\geq\lambda[(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda)]+v_{2}(\lambda,\tilde{e})-\overline{v}_{2}(\lambda)\quad\text{ for }e,\tilde{e}\in E.$
Then we have the following inequality and see that (5) holds.
$\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)$
$=\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]}$
$\geq\lambda[(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda)]+v_{2}(\lambda,\tilde{e})-\overline{v}_{2}(\lambda)$
$=\lambda\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\big{]}\quad\text{ for }e,\tilde{e}\in E.$
On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (6) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the government must find the closest utility control variable for which there is a certainty equivalent to the promised level of lifetime utility. That is, for given $\lambda$, if there is $\hat{e}\in E$ such that $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\hat{e})-\overline{v}_{2}(\lambda))\geq S$, for reported income $e\in E$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},\overline{e}_{\lambda}\})])\big{)},$ (7)
where $\overline{e}_{\lambda}$ is defined such that
$\overline{e}_{\lambda}=\arg\max_{e^{\prime}\in E_{N}}\big{(}v_{2}(\lambda,e^{\prime})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}$
subject to
$\overline{v}_{1}(\lambda)+\frac{1}{\lambda\beta}\big{(}v_{2}(\lambda,e^{\prime})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}<S,$
where $E_{N}\supset E$ is a set with $|E_{N}|=N>M$, $N\in\mathbb{N}$, and for $e\in E_{N}$ $e^{1}\leq e\leq e^{M}$. [^2] To avoid getting stuck on a pass where the same utility control variable $\lambda$ goes on forever, if you find $\overline{e}_{\lambda}=e^{1}$ for given $N$, choose $N^{\prime}\in\mathbb{N}$ with $N^{\prime}>N$, and use $E_{N^{\prime}}$ instead of $E_{N}$ so as to obtain $\overline{e}_{\lambda}\geq e^{2}$. This adjustment also changes the transfers in the current period $t$. However, we see the recursive relation in (7) satisfies the following corresponding incentive constraint.
$(1-\beta)u(e$ $+\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))$
$\geq(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda}\}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\text{ for }e,\tilde{e}\in E.$ (8)
This is because the transfer $\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda)$ in (8) is the solution to the problem (1) subject to $\tau(e_{t})\geq(u^{\prime})^{-1}(\lambda^{-1})-\overline{e}_{\lambda}$ for $t\in\mathbb{N}\cup\{0\}$, where $\lambda=\lambda_{0}$. In other words, since the transfer to the government is constant for individuals whose income is above the threshold $\overline{e}_{\lambda}$, they are indifferent between reporting and misreporting their true income.
With respect to the promise-keeping constraint, it holds because the expectation of the left-hand side of the inequality in (8) is greater than or equal to the promised level of lifetime utility $\overline{v}_{1}(\lambda)$:
$\mathrm{E}\big{[}(1-\beta)u(e+\tau^{*}(\min(e,\overline{e}_{\lambda})|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime})\big{]}\geq\overline{v}_{1}(\lambda).$
Since the definition of $\lambda^{\prime}$ in (7) is an extension of (6), we can use it for the case where $\lambda^{\prime}$ could also be defined in (6). We define a mechanism like this.
Definition 3.
An inter-period transfer mechanism $\mathcal{M}$ is represented by a sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ which satisfies (7) and the corresponding transfers, $\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$.
A sufficient condition for the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (6) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal transfer $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$. Write $F(\lambda)=\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))$ for $\lambda>0$.
Lemma 1.
The sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (6) if
$-\frac{u^{\prime\prime}((u^{\prime})^{-1}(\lambda^{-1}))}{u^{\prime}((u^{\prime})^{-1}(\lambda^{-1}))}<\frac{\beta}{(1-\beta)(e^{M}-\mathrm{E}[e_{t}])}\quad\text{for}\quad\lambda>0.$ (9)
Proof.
It suffices to show the proposition that $F(\lambda)<S$ for all $\lambda>0$.
Since the utility function $u$ is continuous, bounded from above and satisfies the Inada condition, we have
$\lim_{\lambda\to\infty}\overline{v}_{1}(\lambda)=u(\lim_{\lambda\to\infty}(u^{\prime})^{-1}(\lambda^{-1}))\leq S$
$\lim_{\lambda\to\infty}(u^{\prime})^{-1}(\lambda^{-1})=\infty.$
So we see that
$\lim_{\lambda\to\infty}F(\lambda)=\lim_{\lambda\to\infty}\overline{v}_{1}(\lambda)=S.$
If (9) holds, then we have $F^{\prime}(\lambda)>0$ for $\lambda>0$, which implies that $F$ is strictly increasing in $\lambda$. Consequently, we see the desired result $F(\lambda)<S$ for all $\lambda>0$. ∎
If condition (9) does not hold, there may be a case where an utility control variable $\lambda_{t}$ in $\mathcal{M}$ is defined in (7) instead of in (6), e.g. if the individuals’ utilities have constant absolute risk aversion, and if it is greater than $\beta/((1-\beta)(e^{M}-\mathrm{E}[e_{t}]))$. Since $F$ decreases in $\lambda$ and $\lim_{\lambda\to 0}F(\lambda)=\infty$, there must be a threshold $\overline{\lambda}$ below which $F(\lambda)\geq S$.
The inter-period transfer mechanism $\mathcal{M}$ is sequentially incentive compatible regardless of whether each $\lambda_{t}$ is defined in (6) or in (7). Furthermore, if $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (6), like the $\lambda$ mechanism of Marcet and Marimon [1992], $\mathcal{M}$ is sequentially efficient.
Proposition 1.
$\mathcal{M}$ is sequentially incentive compatible. Furthermore, if condition (9) holds, then $\mathcal{M}$ is sequentially efficient.
Proof.
With respect to $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ defined throughout in (6), the propositional statement is shown by Marcet and Marimon [1992]. For completeness, we report the entire proof.
We first show that $\mathcal{M}$ is sequentially incentive compatible. For $\lambda_{t}>0$ with $F(\lambda_{t})<S$
$\lambda_{t}\big{[}$ $(1-\beta)u(e+\tau^{*}(e|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e|\lambda_{t}))\big{]}$
=(6)${=}\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e)-\overline{v}_{2}(\lambda_{t})$ )
$\geq\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},\tilde{e})-\overline{v}_{2}(\lambda_{t})$
$=\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(\tilde{e}|\lambda_{t}))\big{]}\text{ for }e,\tilde{e}\in E.$
The last inequality follows from the optimality of $\lambda\overline{v}_{1}(\lambda)+v_{2}(\lambda,e)$ in the problem (1) given $\lambda$. Hence, $\mathcal{M}$ is sequentially incentive compatible for $\lambda>0$ defined in (6).
For $\lambda_{t}>0$ with $F(\lambda_{t})\geq S$
$\lambda_{t}\big{[}(1-\beta)$ $u(e+\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t}^{\prime}(e|\lambda_{t}))\big{]}$
=(7)${=}\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}$
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{e,\overline{e}_{\lambda_{t}}\})-\mathrm{E}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
$\geq\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}$
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{\tilde{e},\overline{e}_{\lambda_{t}}\})-\mathrm{E}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
$=\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t}^{\prime}(\tilde{e}|\lambda_{t}))\big{]}\text{ for }e,\tilde{e}\in E.$
The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
We proceed to show that if condition (9) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms. By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible. By Lemma 1 since condition (9) holds, the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (6). Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period. Therefore $\mathcal{M}$ is a Pareto optimal transfer mechanism.
It remains to prove that under condition (9) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism. Suppose, contrary to our claim, that there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$. Let $(v_{1}^{*},v_{2}^{*})$ be the present value achieved by $\Gamma$. Set $\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*})$. Since condition (9) holds for any $\lambda>0$, this $\lambda_{0}$ satisfies $F(\lambda_{0})<S$ as in the proof of Lemma 1. We may now use $(\lambda_{0},e)$ as the initial condition for $\mathcal{M}$ under condition (9). Then, by construction, each individual has the same present value for both contracts. Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$. However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under condition (9). ∎
Proposition 1 shows sequential efficiency of the mechanism $\mathcal{M}$ only for the case where condition (9) holds. However, we could relax this restriction by removing the assumption that the utility function is bounded from above. This assumption provides a sufficient condition to ensure Proposition 2, but it is not a necessary condition.
3 The inter-period transfer mechanism for each individual
The government’s approach of using the mechanism $\mathcal{M}$ successfully avoids an immiserization outcome. That is, the promised lifetime utility of each individual in $\mathcal{M}$ converges almost everywhere to a random variable with a finite expectation. To see this, we reinterpret $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}}$ as a sequence of random variables on $(\Omega,\mathcal{F},P)$. This is done recursively as follows. First we interpret $\overline{v}_{1}(\lambda_{1})$ as a random variable $\overline{v}_{1}(\lambda_{1}(e_{0}(\cdot)|\lambda_{0}))$ and let $\mathcal{F}_{1}\subset\mathcal{F}$ be the smallest $\sigma$-algebra induced by $\overline{v}_{1}(\lambda_{1})$. That is, $\mathcal{F}_{1}=\{\overline{v}_{1}(\lambda_{1})^{-1}(B)\ |\ B\in\mathcal{B}(\mathbb{R})\}$, where $\overline{v}_{1}(\lambda_{1})^{-1}(B)=\{\omega\in\Omega\ |\ \overline{v}_{1}(\lambda_{1}(e_{0}(\omega)|\lambda_{0}))\in B\}$. We see that $\overline{v}_{1}(\lambda_{1})$ is $\mathcal{F}_{1}$-measurable, $\overline{v}_{1}(\lambda_{1}):(\Omega,\mathcal{F}_{1})\to(\mathbb{R},\mathcal{B}(\mathbb{R})).$ Second, we consider $\overline{v}_{1}(\lambda_{n})$ as a random variable $\overline{v}_{1}(\lambda_{n}(e_{n-1}(\cdot)|\lambda_{n-1}))$ and define $\mathcal{F}_{n}\subset\mathcal{F}$ as the smallest $\sigma$-algebra induced by the product of the elements in $\{\overline{v}_{1}(\lambda_{i})\}^{n}_{i=1}$ such that
$\mathcal{F}_{n}=\{(\overline{v}_{1}(\lambda_{1})\times\dots\times\overline{v}_{1}(\lambda_{n}))^{-1}(B)\ |\ B\in\mathcal{B}(\mathbb{R}^{n})\},$
where
$(\overline{v}_{1}(\lambda_{1})\times\dots\times$ $\overline{v}_{1}(\lambda_{n}))^{-1}(B)$
=$=\Big{\{}\omega\in\Omega\ \Big{|}\ \big{(}\overline{v}_{1}(\lambda_{1}(e_{0}(\omega)|\lambda_{0})),\dots,\overline{v}_{1}(\lambda_{n}(e_{n-1}(\omega)|\lambda_{n-1}))\big{)}\in B\Big{\}}.$
We see that $\overline{v}_{1}(\lambda_{n})$ is $\mathcal{F}_{n}$-measurable, $\overline{v}_{1}(\lambda_{n}):(\Omega,\mathcal{F}_{n})\to(\mathbb{R},\mathcal{B}(\mathbb{R})).$
Proposition 2.
The sequence of promised lifetime utilities of individuals $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ converges to an integrable random variable.
Proof.
We first show that $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale. Since the sequence $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}\in\mathcal{M}$ satisfies (7), we have
$\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})(\omega)]=\overline{v}_{1}(\lambda_{n-1})(\omega)$
for given realised $\overline{v}_{1}(\lambda_{n-1})(\omega)$. We thus have the following equalities: For all $B\in\mathcal{B}(\mathbb{R})$ for all $n\in\mathbb{N}$
[…]
and
[…]
From the last equality we have for $n\in\mathbb{N}$ $\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e. $[P]$, and $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale as claimed.
We proceed to show that $\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$. From (6) and (7) we have $\overline{v}_{1}(\lambda_{n})\leq S$ for $n\in\mathbb{N}$. So we have $\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$.
Since $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale and $\sup_{n}\mathrm{E}[\overline{v}_{1}(\lambda_{n})^{+}]\leq\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$, from the submartingale convergence theorem, there is an integrable random variable $v^{\infty}$ such that $\overline{v}_{1}(\lambda_{n})\to v^{\infty}$ almost everywhere. ∎
As the proof of Proposition 2 shows, mechanism $\mathcal{M}$ functions like a fair-value gamble on lifetime utilities across periods. The lifetime utility in each period is a random variable whose sequence forms a martingale. Consequently, the mechanism is fair in the sense that each individual’s expected lifetime utility remains unchanged from the time of the bet.
4 The inter-period transfer mechanism for the government
However, the convergence of individuals’ lifetime utilities to an integrable random variable with finite expectation does not yet ensure the sustainability of the $\mathcal{M}$ mechanism. This is because the government predicts a declining trend in its budget for each individual in the long run, as we see in the following lemma. Let $\overline{\tau}_{n}$ be the average transfer to those who were born $n$ periods ago. Then $\overline{\tau}_{n}$ is given by as follows.
$\overline{\tau}_{n}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathrm{E}[e_{n}]\quad\text{ for }n\in\mathbb{N}\cup\{0\}.$
Lemma 2.
The average transfer for each age of individuals $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is an increasing sequence.
Proof.
We first consider the case $n=1$. For $n=1$, since $\mathrm{E}[\overline{v}_{1}(\lambda_{1}(e_{1}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
$\overline{\tau}_{1}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathrm{E}[e_{1}]$
$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathrm{E}[e_{0}]=\overline{\tau}_{0}.$
For $n>1$ since $u^{-1}$ is a convex function, from Jensen’s inequality we see the result.
$\overline{\tau}_{n-1}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n-1}))]-\mathrm{E}[e_{n-1}]$
$=\mathrm{E}[u^{-1}(\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})])]-\mathrm{E}[e_{n-1}]$
$\leq\mathrm{E}[\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))|\overline{v}_{1}(\lambda_{n-1})]]-\mathrm{E}[e_{n-1}]$
$=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathrm{E}[e_{n}]$
$=\overline{\tau}_{n}\quad\text{ for }n=2,3,\cdots.$
The proof of Lemma 2 shows that the increasing trend in $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is due to the cost of keeping the expected lifetime utility in each period unchanged from the previous period, although there are risks between periods.
Let $g(t,\lambda_{0})$ be the expected balance of payments in period $t$ for the government for given initial utility control variable $\lambda_{0}$. Since each individual’s probability of being alive is $\alpha\in(0,1)$, and from period one the next generation of contimuum of individuals on the $(1-\alpha)$ interval is born in each period, $g(t,\lambda_{0})$ is given by
$g(t,\lambda_{0})$ $=\alpha^{t}(-\overline{\tau}_{t})+\alpha^{t-1}(1-\alpha)(-\overline{\tau}_{t-1})+\cdots+\alpha(1-\alpha)(-\overline{\tau}_{1})+(1-\alpha)(-\overline{\tau}_{0})$
$=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}.$
For the $\mathcal{M}$ mechanism to be sustainable, the balance of payments in each period must be greater than or equal to be zero. As we see in the following proposition, with the approximation of the second order Taylor polynomial, if incomes have a symmetric distribution, a sufficient condition for the balance of payments in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
Proposition 3.
Assume that (i) incomes have a symmetric distribution, (ii) $-u^{\prime\prime}/u^{\prime}$ is a decreasing function, (iii) $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function or equivalently,
$\frac{\left(-\frac{u^{\prime\prime}}{u^{\prime}}\right)^{{}^{\prime\prime}}({\scriptstyle u^{-1}(y)})}{-\left(-\frac{u^{\prime\prime}}{u^{\prime}}\right)^{\prime}({\scriptstyle u^{-1}(y)})}\leq\frac{(u^{-1})^{\prime\prime}(y)}{(u^{-1})^{\prime}(y)},$
and (iv) $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. Then $g(t,\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ with the approximation of the second order Taylor polynomial.
The proof of Proposition 3 uses the following lemma.
Lemma 3.
Under the assumptions (i), (ii), (iii) in Proposition 3, the second order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
Proof.
First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (6). Using the third order Taylor polynomial, for a given utility control variable $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
$u^{-1}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=$ $u^{-1}(\overline{v}_{1}(\lambda_{n-1}))+\frac{1}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\Delta_{n-1}-\frac{1}{2!}\frac{u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{3}}\Delta_{n-1}^{2}$
$+\frac{1}{3!}\left(-\frac{u^{\prime\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{4}}+\frac{3u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{2}}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{5}}\right)\Delta_{n-1}^{3}$
$+h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))\Delta_{n-1}^{3},$ (10)
where $\Delta_{n-1}=(1-\beta)(e-\mathrm{E}[e_{n}])/(\beta\lambda_{n-1})$ and $h_{3}:\mathbb{R}\to\mathbb{R}$ is a function such that $\lim_{\Delta_{n-1}\to 0}h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=0$. Taking the expectation of (10), $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is written as
$\overline{\tau}_{n}-\overline{\tau}_{n-1}$ $=\mathrm{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}[e_{n}]\right]+o(\Delta_{n-1}^{3}).$
Note that since incomes follow a symmetric distribution, the third term on the right-hand side of the equation (10) will cancel out when the expectation is taken.
Since $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function, from Jensen’s inequality we have
$\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\left(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))\right)\right]$ $=\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}\big{(}\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})]\big{)}\big{)}\right]$
$\geq\mathrm{E}\left[\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n}))\big{)}\bigg{|}\overline{v}_{1}(\lambda_{n-1})\right]\right]$
$=\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
It follows that the third order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence.
We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (7). We change $\Delta_{n-1}$ in (10) to
$\overline{\Delta}_{n-1}=\frac{(1-\beta)\big{(}\min\{e,\overline{e}_{\lambda_{n-1}}\}-\mathrm{E}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\big{)}}{\beta\lambda_{n-1}}.$
By the asymmetric distribution of conditional income $e\leq\overline{e}_{\lambda_{n-1}}$, the third term on the right-hand side of (10) remains and $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is given by
$\overline{\tau}_{n}-\overline{\tau}_{n-1}=\mathrm{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\right]+o(\overline{\Delta}_{n-1}^{2}).$
By a similar argument, the second order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
We prove Proposition 3 below.
Proof.
From Lemma 3, with the second order polynomial approximation we have
$\overline{\tau}_{n}-\overline{\tau}_{n-1}\leq\overline{\tau}_{1}-\overline{\tau}_{0}\quad\text{for }n\in\mathbb{N}.$
Hence the following inequality holds.
$g(t,\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
$\geq\alpha\cdot\frac{1-\alpha^{t}}{1-\alpha}(\overline{\tau}_{0}-\overline{\tau}_{1})-\overline{\tau}_{0}.$
The last term is positive if
$-\overline{\tau}_{0}\geq\frac{\alpha(1-\alpha^{t})}{1-\alpha}(\overline{\tau}_{1}-\overline{\tau}_{0})$
and we see a sufficient condition for $g(t,\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ is $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Proposition 3 is given by the third order polynomial approximation as follows.
Corollary 1.
The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Proposition 3 holds with the second order polynomial approximation if the initial utility control variable $\lambda_{0}$ satisfies
$\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))\geq\sqrt{\frac{(1-\alpha r)^{2}}{\alpha r^{2}(1-\alpha)}\mathrm{Var}[e_{t}]},$ (11)
where $\lambda_{E}=1/(u^{\prime}(u(\mathrm{E}[e_{t}])))$.
Proof.
Let us first examine the case where $\lambda_{1}$ can be defined in (6). Applying Tayler theorem to $u^{-1}(\overline{v}_{1}(\lambda_{1}))$, $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)$ is expressed as
$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{0}]+o(\Delta_{0}^{3}).$ (12)
On the other hand, applying Tayler theorem to $u^{-1}(\overline{v}_{1}(\lambda_{E}))$, we have
$-\overline{\tau}_{0}$ $=-u^{-1}(\overline{v}_{1}(\lambda_{0}))+u^{-1}(\overline{v}_{1}(\lambda_{E}))$
$=\frac{(u^{-1})^{\prime}(\overline{v}_{1}(\lambda_{0}))}{1!}\Delta_{E}+\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\Delta_{E}^{2}+o(\Delta_{E}^{2}),$ (13)
where $\Delta_{E}=\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0})$. Since the first term on the right-hand side of (13) is positive, it is sufficient to show the second term on the right-hand side of (13) is equal to or greater than the first term on the righ-hand side of (12). However, from the corollary assumption (11), we have
$\Delta_{E}^{2}\geq\frac{(1-\beta)^{2}}{\lambda_{0}^{2}\beta^{2}}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{t}],$
which is the desired conclusion that the condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ holds with the second order polynomial approximation. For the case where $\lambda_{1}$ must be defined in (7), we have
$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}\big{[}\min\{e_{0},\overline{e}_{\lambda_{0}}\}\big{]}+o(\overline{\Delta}_{0}^{2}).$
Since $\mathrm{Var}[\min\{e_{0},\overline{e}_{\lambda_{0}}\}]\leq\mathrm{Var}[e_{0}]$, (11) is sufficient to be $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
Figure 1: Slope of Utility Function
{singlespace*}
Notes: This figure shows a graph of the slope of a utility function. The horizontal axis is consumption, and the vertical axis is utility. The curve is the utility function. The length $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$ is a linear approximation of the length between the expected income and the consumption guaranteed in the mechanism M with $\lambda_{0}$.
The left-hand side of the inequality (11) is a linear approximation of the difference between the consumption level that ensures the expected income and the consumption level in the mechanism $\mathcal{M}$ with the utility control variable $\lambda_{0}$. The inequality (11) means that this length is greater than the square root of the weighted income variance. Figure 1 shows which part of the graph of a utility function corresponds to $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$. The lifetime utility $\overline{v}_{1}(\lambda_{0})$ in the mechanism $\mathcal{M}$ could be set higher than that in the state of autarky if the risk premium of the state of autarky is higher than the term in the right-hand side of the inequality (11).
A higher discount factor and higher probability of being alive will relax the sufficient condition of $\lambda_{0}$. If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (11) may be unnecessarily strict.
5 Numerical examples
According to Corollary 1, in the following case, the mechanism $\mathcal{M}$ is sustainable and the lifetime utilities of all generations are higher than that in the state of autarky. Suppose that the utility function is given by
$u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma},$
where $\gamma=-3$. The income set $E$ is $E=\{0,1,2,3,4,5,6,7,8,9\}$. We will refer to this income set as $E_{s}$. The income shocks follow the distribution $P(e_{t}=0)=Pr(e_{t}=9)=0.025$, $P(e_{t}=1)=P(e_{t}=8)=0.045781$, $P(e_{t}=2)=P(e_{t}=7)=0.092762$, $P(e_{t}=3)=P(e_{t}=6)=0.148524$, $P(e_{t}=4)=Pr(e_{t}=5)=0.187933$. We will refer to this distribution as $P_{s}$ Suppose that the discount factor and the probability of being alive are $r=\alpha=0.93$. Then Proposition 3 assumptions are satisfied for $\lambda_{0}\in[\lambda_{CE},133]$, where $\lambda_{CE}=49.608=1/u^{\prime}(E[u(e_{t})])$. Note that $\lambda_{E}=166.375$, the risk premium is $1.826$, and $\sqrt{(1-\alpha r)^{2}/(\alpha r^{2}(1-\alpha))\mathrm{Var}[e_{t}]}$ is $1.147$.
Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals and corresponding transfer balances for $100$ periods for the above case. The initial utility control variable is set $\lambda_{0}=50$.
As Proposition 2 implies, the series of sample means of lifetime utilities does not approach negative infinity, which has been called immiserization. However, as Lemma 2 shows, the corresponding series of sample means of transfer balances $\{-\tau_{n}\}_{n=0}^{100}$ is decreasing.
Figure 2: A series of sample means of lifetime utility and corresponding transfer balance
{singlespace*}
Notes: This figure shows a series of sample means of lifetime utilities and corresponding transfer balances for $100$ periods, where $1000$ individuals are assumed to have the utility $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$. The parameters are set so that the initial utility control variable $\lambda_{0}=50$, the discount factor $r=0.93$, the probability of being alive $\alpha=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated using the discrete probability distribution $P_{s}$.
Figure 3 shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods. In this society, each individual will be alive in the next period with probalibility $\alpha=0.93$, and newly born individuals enter the society, holding the total population $100$. As in Figure 2, each individual’s utility is $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial utility control variable $\lambda_{0}=50$ and the discount factor $r=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated with the same symmetric distribution $P_{s}$.
The series of sample means of transfer balances $\{g(n,\lambda_{0})\}_{n=0}^{1000}$ remains non-negative, and supports Proposition 3.
Figure 3: A series of sample means of balance of payments $\{g(t,\lambda_{0})\}$
{singlespace*}
Notes: This figure shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods. As in Figure 2, each individual’s utility is $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial utility control variable $\lambda_{0}=50$, the discount factor $r=0.93$, the probability of being alive $\alpha=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated using the discrete probability distribution $P_{s}$.
Immiserization observed in previous studies [e.g. Green, 1987, Thomas and Worrall, 1990] is due to the approach to allocate risks within each period across the polulation of individuals. To ensure truthful reporting, only individuals with the highest incomes keep their lifetime utility level in the next period, and lifetime utilities of other individulas in the next period are discounted according to reported incomes. As a result, the average contract value of individuals becomes a decreasing series.
However, the mechanism $\mathcal{M}$ simply shifts risks to future periods. A similar limitation is that, the proposed mechanism leads to a dicreasing series of the average transfer balances for each generation, which could make the government’s balance of payments negative infinity. However, this could be solved by setting the initial utility control variable appropriately, so that average transfers are negative for the first few periods from birth. If this sense of intergenerational cooperation is properly organised, efficient income allocations could be sustainable even under asymmetric information about income shocks.
The property of the mechanism $\mathcal{M}$ contrasts with the numerical results of Marcet and Marimon [1992], where the lifetime utility of agent 1 (manager) is higher in the state of autarky than in the state of the incentive contract with agent 2 (investor). This is probably because, first, Marcet and Marimon [1992] is a growth model with the assumption that capital and the manager’s consumption, increases when the manager chooses the optimal level of investment. Second, since the transfers in the numerical example for incentive contracts in Marcet and Marimon [1992] are not competitive, the external financing contracts end up reducing agent 1’s allocation.
6 Checking the mecanism optimality for each generation
Lastly, we express the mechanism $\mathcal{M}$ using a Bellman equation. We focus on the case that (9) in Lemma 1. First we guess the value function $V:\mathbb{R}_{++}\times\mathbb{N}\cup\{0\}\to\mathbb{R}\cup\{-\infty\}$ is given by for $\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$
$V(\lambda_{t},t)=\mathrm{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}.$ (14)
Second, we verify whether this function satisfies the Bellman equation:
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1),$ (15)
where $\Lambda:\mathbb{R}_{++}\times\mathbb{N}\cup\{0\}\multimap\mathbb{U}$ is an available control correspondence defined such that
$\Lambda(\lambda_{t},t)=\arg\max_{\tau}\left[\lambda_{t}u(e_{t}+\tau)-\tau\ \text{ subject to }(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k}\tau(e_{k})\right]\leq\text{Const.}\right]$. ]
$\mathbb{U}$ is the set of control parameters, and $\lambda_{t+1}$ in (15) is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{U}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (7). We observe the following proposition.
Proposition 4.
If the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ satisfies (9), then $V$ fullfills the Bellman equation and it is the value function of the dynamic optmisation problem, where the control function $\tau^{*}$ maximises $J$ in (2).
A sufficient condition for $V$ to be the value function of the Bellman equation is given by Wiszniewska-Matyszkiel [2011] which is to satisfy the Bellman equation and the following terminal condition.
Terminal condition [Wiszniewska-Matyszkiel, 2011]
  • (i)
    For every $\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$
    $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}\leq 0$
    and
  • (ii)
    for every $\lambda_{t}\in\{\lambda\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$, if $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}<0$, then
    $J(\lambda_{t},t,\hat{\tau})=-\infty$
    for every $\hat{\tau}:\mathbb{R}_{++}\to\mathbb{R}_{+}$ such that $\lambda_{t}$ in $\mathcal{M}$ corresponding to $\hat{\tau}$ and $t$.
Proof.
We observe first that if (9) holds, then $V$ defined in (14) satisfies the Bellman equation in (15). Indeed, since by definition of $\tau^{*}$ we have $u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))=\overline{v}_{1}(\lambda_{t})$,
$V(\lambda_{t},t)$ $=\mathrm{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathrm{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1})|\overline{v}_{1}(\lambda_{t})\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1).$ (16)
The last equation follows from the fact that $\lambda_{t+1}$ is realized given $e_{t}$ and seeing $\mathrm{E}_{t+1}$ as a function of $e_{t+1}$. But $\tau^{*}$ is the solution to the problem
$\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}))+\beta\mathrm{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1})|\overline{v}_{1}(\lambda_{t})\big{]}$ (17)
under the binding constraint
$(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k}\tau(e_{k})\right]=\mathrm{E}[\tau(e_{t})]=\text{Const.}$. (18)
Indeed, for a given $\lambda_{t}$ the objective function in (17) multiplied by $\lambda_{t}$ is written as
$\lambda_{t}\big{[}(1-\beta)$ $u(e_{t}+\tau(e_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\mathrm{E}[v_{2}(\lambda_{t},e_{t})]$
$=(1-\beta)\big{[}\lambda_{t}u(e_{t}+\tau(e_{t}))-\tau(e_{t})\big{]}+\beta\lambda_{t}\overline{v}_{1}(\lambda_{t})+(1-\beta)\mathrm{E}[\tau(e_{t})],$
and we see that the last terms are maximized by $\tau^{*}$ under the constraint (18). It follows that (16) is written as
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1),$
and we see the desired result.
The task is now to check the terminal condition. Since $u$ is bounded from above, we have
$\lim_{t\to\infty}$ $V(\lambda_{t},t)\beta^{t}=0.$
Thus, $V$ satisfies Terminal condition (i) and from Theorem 1 of Wiszniewska-Matyszkiel [2011], we see $V$ defined in (14) is the value function of the Bellman equation in (15), and the control function $\tau^{*}$ maximises $J$ in (2). ∎
7 Conclusion
This paper presents an alternative solution to avoid the prediction of the previous work: if the incomes of infinitely lived individuals are unobservable, efficient allocations are achieved only at the cost of invoking permanent inequality, leading to an immiserization of society. The proposed inter-period transfer mechanism achieves within-period full insurance by postponing risks. It does not trade off equal opportunities to become wealthy. It could be sustained by intergenerational cooporation. The result sheds light on efficient resource allocation for sustainable societies with equal opportunities.

### v15

@@
-
+Insuring uninsurable income
 Abstract
-In an exchange economy composed of risk-averse individuals facing shocks, these shocks may become uninsurable if they are private information.
-Efficient shock smoothing across the economy requires self-reporting of the shocks that each individual privately suffers.
-This efficient allocation of resources leads to growing inequality between individuals, and the problem remains largely unsolved.
-This paper proposes a solution.
-The proposed mechanism (i) promises within-period full insurance by postponing risks ; (ii) does not lead to social ranking among the same age groups; and (iii) is sustainable under some conditions.
-Key words: Incentives, optimal resource allocation, recursive contract, equality of same age group.
+We study dynamic mechanism design in a pure-exchange economy with privately observed idiosyncratic income.
+Classic hidden-income contracts attain constrained efficiency only at the cost of immiseration (Green 1987; Thomas–Worrall 1990).
+We propose a simple recursive mechanism—adapted from Marcet–Marimon (1992)—that shifts each income shock forward by one period, keeps promised utilities in a bounded set, and, under a transparent “moderate risk-aversion” condition, delivers sequential efficiency.
+In a stationary overlapping-generations setting, we further provide an explicit condition on the initial promise that ensures budget sustainability; early cohorts pre-fund intertemporal smoothing so that every cohort attains a higher expected lifetime utility than under autarky.
+Our analysis uses a single state (promised utility), closed-form transfers, and a Bellman verification.
+Keywords: Dynamic mechanism design; recursive contracts; private information; immiseration; overlapping generations.
+JEL: D82, D86.
 1 Introduction
-In an exchange economy composed of risk-averse individuals facing shocks, these shocks may become uninsurable if they are private information.
-This problem has been studied extensively since the work of Green [1987].
-These shocks are uninsurable because efficient shock smoothing across the economy requires self reporting of shocks that each individual privately suffers.
-Proposed solutions are limited.
-Inducing truth-telling at the expense of equality leads to immiserization, where the average expected lifetime utility of individuals converge to negative infinity while driving inequality without bound[Green, 1987, Thomas and Worrall, 1990].
-If discount factor is approximately one, meaning individuals value the future and care little about the present, their contracts approach to the first best [Thomas and Worrall, 1990, Carrasco et al., 2019].
-Atkeson and Lucas [1992] consider a scenario where discount factors represent resource prices, and social planners trade claims to current and future resources at these prices.
-They show that full information is required for efficient allocation.
-This paper extends an alternative mechanism proposed by Marcet and Marimon [1992], which distributes risk across periods for the same individual.
-This approach could reduce the inequality observed in previous studies that allocate risk among individuals within each period.
-The mechanism proposed by Marcet and Marimon [1992] was an efficient growth mechanism combining state-contingent investment and transfers.
-This paper uses it for transfers in an exchange economy facing income shocks.
-The first key finding of this paper is that our mechanism can achieve both efficiency and avoidance of immiserization under certain conditions   specifically, when individuals’ degree of risk aversion is not excessively high relative to the disparity between their highest and lowest possible incomes.
-Furthermore, the mechanism preserves the equality of opportunity within the same age group.
-Both efficiency and equality of opportunity are achieved by shifting risks to future periods.
-However, if the degree of risk aversion is too high relative to the disparity in income levels, challenges may arise.
-In such cases, it may become necessary to temporarily adjust the optimal transfer rules.
-This disruption occurs when the promised lifetime utility cannot be defined, given the assumption that each individual’s utility function is bounded from above.
-Nevertheless, even when the sufficient condition is not met, the mechanism continues to maintain the optimal allocation under the given condition and avoids immiserization.
-The second key finding pertains to the potential sustainability of this mechanism in a society with a continuum of individuals, where the population, consisting of different age group, remains constant over time.
-In such a society, each individual faces a probability of dying in the next period, while new individuals continuously enter.
-Since risks exist across periods, maintaining unchanged expected lifetime utility across periods incurs costs.
-To ensure sustainability, earlier entrants must bear these costs, resulting in average consumption increasing with age.
-However, this does not lead to the formation of social ranking within the same age group.
-Importantly, this mechanism can ensure higher expected lifetime utility compared to autarky for all generations, while sustaining a positive balance of payments over time.
-Thus, a society that achieves both efficiency and equality of opportunity is feasible.
-The paper is structured as follows.
-Section 2 introduces the notation and tarminology and proposes the mechanism for the inter-period risk transfers.
-Section 3 shows efficiency of the proposed mechanism and examines the properties for individuals.
-Section 4 investigates the potential sustainability of the mechanism.
-Section 5 provides some numerical examples.
-Section 6 evaluates the mechanism’s efficiency again using Bellman equation.
-Section 7 summarizes the conclusions.
+Private information can render idiosyncratic income effectively uninsurable.
+Since the seminal work of Green [1987] and Thomas and Worrall [1990], the constrained-efficient hidden-income contract trades insurance off against incentives and features a drift of average promised utilities toward a lower bound—immiseration—with cross-sectional dispersion rising over time.
+[^2] Earlier papers sometimes use immiserization[e.g., Phelan, 1998, Zhang, 2009].
+Recent work with persistent private information, Bloedel et al.
+[2025], also sharpens the mechanism of immiseration and explains the balance of the cost of incentive provision and backloading of high-powered incentives.
+When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but the knife-edge offers little comfort for moderate discounting.
+A different route was proposed by Marcet and Marimon [1992].
+Allowing agents to make productivity-enhancing investments, they designed a transfer rule that shifts risk across periods for the same individual.
+Capital accumulation enlarges the future consumption set and breaks the immiseration trap; the associated recursive saddle-point representation, later formalized by Marcet and Marimon [2019] makes the analysis more tractable.
+Yet the mechanism has not been explored in a pure-exchange economy without investment or storage, where the scope for avoiding downward drift is a priori unclear.
+This paper adapts the Marcet–Marimon risk-shifting mechanism to an exchange economy in which agents cannot save.
+Promised utility is the sole state variable, so the model yields a closed-form characterization.
+If individual assets were allowed, one would require the set-valued approximation methods of Sleet and Yeltekin [2016]—a contrast confirmed by our closed-form example in Section 4.
+The contribution of my paper is threefold.
+(i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohort pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts.
+The organization of the paper is as follows.
+Section 2 sets up the model and transfer rule.
+Section 3 proves efficiency and boundedness, and analyzes inter-generational sustainability.
+Section 4 presents numerical illustrations.
+Section 5 verifies optimality via a Bellman equation, and Section 6 concludes.
 2 Model
-Consider an infinitely existing risk-neutral government for a society with a continuum of individuals in a unit interval.
-Each individual is risk-averse and will be alive in the next period with probability $\alpha\in(0,1)$.
-In each period, newly born risk-averse individuals enter the society, holding the total population constant.
-Individuals differ only in their periods of birth and death.
-Each individual has a discount factor $r\in(0,1)$.
-Let $\beta=\alpha r$ be the multiplication of $r$ and $\alpha$.
-A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game.
-Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $\{e_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ be a sequence of integrable random variables on $(\Omega,\mathcal{F},P)$ denoting identically and independently distributed income shocks of each individual in a sequence of periods.
-These income shocks take values in $E:=\{e^{1},e^{2},\dots,e^{M}\}$, where $e^{i}\in\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$.
-The government considers a transfer mechanism $\mathcal{M}$ that repeats from period zero to infinity, using a sequence of transfers (controls) $\{\tau_{t}\}_{t=0}^{\infty}$, $\tau_{t}\in\mathbb{R}$ and a sequence of utility control variables $\{\lambda_{t}\}_{t=0}^{\infty}$, $\lambda_{t}\in\mathbb{R}_{++}$ which will be specified later.
-2.1 Pareto optimal transfer mechanism with a utility control variable $\lambda_{0}$
-First, we derive an optimal transfer function which, under symmetric income information, ensures a level of utility corresponding to a utility control variable $\lambda_{0}>0$ Following Marcet and Marimon [1992], we will use this utility control variable $\lambda_{0}$ as a state variable under asymmetric income information.
-For this purpose, in this subsection only, we consider the case where individuals cannot hide their income from the government in each period.
-In period 00, the government is considering an efficient transfer mechanism that increases individuals’ lifetime utility as much as possible, but the government cannot allow an infinite net subsidy.
-The government solves the following problem, having decided on an upper bound for the net subsidy.
-()$(\ast)\begin{cases}\quad\max_{\{\tau_{t}\}_{t=0}^{\infty}}(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})\right]\\ \text{subject to}\\ \quad(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}(\tau_{t})\right]\leq\text{Const},\end{cases}$
-where $\mathrm{E}_{0}$ is the conditional expectation given $e_{0}$, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period.
+2.1 Environment
+We consider a risk-neutral planner allocating transfers across a continuum of risk-averse individuals indexed by birth cohorts.
+Individuals survive to the next period with probability $\alpha\in(0,1)$ and discount at rate $r\in(0,1)$; their effective discount factor is $\beta=\alpha r$.
+In each period, a newborn cohort enters so that the cross-section is stationary.
+[^3] A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game.
+Each individual draws an i.i.d.
+income shock $e_{t}\in E:=\{e^{1},e^{2},\dots,e^{M}\}\subset\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$.
+The planner commits to a recursive transfer mechanism $\mathcal{M}$ mapping the current promised $\lambda_{t}>0$ into a transfer $\tau_{t}$ and next period’s promise $\lambda_{t+1}$.
+The single state variable is the promised utility $\overline{v}_{1}(\lambda)$, which is strictly increasing in $\lambda$.
+2.2 Symmetric-information benchmark
+First, we derive an optimal transfer function which, under symmetric income information, ensures a level of utility corresponding to a promise $\lambda_{0}>0$ .
+Following Marcet and Marimon [1992], we will use this promise $\lambda_{0}$ as a state variable under asymmetric income information.
+For this purpose, in this subsection only, we consider the case where individuals cannot hide their income from the planner in each period.
+At period 0, the planner is considering an efficient transfer mechanism that increases individuals’ lifetime utility as much as possible, but it cannot afford an infinite net subsidy.
+The planner solves the following problem, which is normalized by multiplying $(1-\beta)$, having decided on an upper bound for the net subsidy.
+$(\ast)\quad\max_{\{\tau_{t}\}_{t=0}^{\infty}}(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})\right]\quad\text{s.t.
+}\quad(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}(\tau_{t})\right]\leq\overline{C},$
+where $\overline{C}>0$ is an exogenous upper bound on the present value of net subsidies, $\mathbb{E}_{0}$ is the conditional expectation given the information in period 0, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period.
 The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
-Without specifying the upper bound of the net subsidy, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines the upper bound of the net subsidy as well.
-Thus, the problem ()($\ast$)( ∗ ) can be rewritten as the following problem, where the objective function is the Lagrangian of problem ()($\ast$)( ∗ ) with an undetermined multiplier $01/\lambda_{0}>$ 0.
-$\max_{\{\tau_{t}\}_{t=0}^{\infty}}$ $(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}[\lambda_{0}u(c_{t})-\tau_{t}]\right].$ (1)
+Without specifying $\overline{C}$, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines $\overline{C}$ as well.
+Let $\mu>0$ be the multiplier on the subsidy constraint and put $\lambda_{0}=1/\mu$.
+The problem $(\ast)$ can be rewritten as the following problem.
+$\max_{\{\tau_{t}\}_{t=0}^{\infty}}$ $(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}[\lambda_{0}u(c_{t})-\tau_{t}]\right].$ (1)
 From the first order condition,
 $u^{\prime}(c_{t})=\frac{1}{\lambda_{0}}\text{ for }t\in\mathbb{N}\cup\{0\}$
-and we obtain $c_{t}=(u^{\prime})^{-1}(\lambda_{0}^{-1})$ and $\tau_{t}=(u^{\prime})^{-1}(\lambda_{0}^{-1})-e_{t}$ for $t\in\mathbb{N}\cup\{0\}$.
-This is the Pareto optimal transfer mechanism that gives the individual constant consumption in each period.
-Let $v_{1}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing each individual’s lifetime utility provided by the transfer mechanism.
+and the optimal policy delivers a constant consumption $c^{\ast}(\lambda_{0})=(u^{\prime})^{-1}(\lambda_{0}^{-1})$ and transfer $\tau^{\ast}(e|\lambda_{0})=(u^{\prime})^{-1}(\lambda_{0}^{-1})-e$, where $\tau^{*}(\cdot\ |\lambda_{0})$ is a time-invariant transfer function.
+We record two objectives for future use: Let $v_{1}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing each individual’s lifetime utility provided by the transfer mechanism.
 Let $v_{2}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing the expected balance of transfers.
 Then, for given $\lambda_{0}>0$ and $e_{0}\in E$, these values are written as follows.
-$v_{1}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathrm{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})]=u((u^{\prime})^{-1}(\lambda_{0}^{-1}))=:\overline{v}_{1}(\lambda_{0})$
-$v_{2}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathrm{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}(-\tau_{t})].$
-Note that the function $v_{1}$ could be seen as a function of the value $\lambda_{0}$ whose inverse was the Lagrangian multiplier.
-We write it as $\overline{v}_{1}$.
-The function $\overline{v}_{1}$ is bijective and increasing.
-On the other hand, we see that $v_{2}(\lambda_{0},e_{0})$ depends on income.
-So we write its expectation as $\overline{v}_{2}(\lambda_{0})$.
-$\overline{v}_{2}(\lambda_{0})$ $:=\mathrm{E}[v_{2}(\lambda_{0},e_{t})].$
-In Section 4 the government will adjust the value of $\lambda_{0}$ so that the expected balance of total transfers across generations is non-negative.
-Let $\tau^{*}(\cdot\ |\lambda_{0})$ be a time-invariant transfer function, which is the solution to the problem (1).
-That is, $\tau^{*}:\mathbb{R}_{+}\times\mathbb{R}_{++}\to\mathbb{R}_{+}$ is defined such that, for a reported income $x\in\mathbb{R}_{+}$ and a utility control variable $\lambda>0$
-$\tau^{*}(x|\lambda):=(u^{\prime})^{-1}(\lambda^{-1})-x.$
-We will use these notation in the following section for the case where income is private information.
-There we observe that a utility control variable $\lambda_{t}$ can play the role of a state variable and the role of an incentive device.
-Let $\mathcal{T}$ be the set of all admissible control functions there, and let $\tau:\mathbb{R}_{+}\times\mathbb{R}_{++}\to\mathbb{R}_{+}$ be a typical element of $\mathcal{T}$.
-Then the lifetime utility of individuals who are guaranteed utility level $\lambda_{t}$ at an initial period $t$ is given by
-$J(\lambda_{t},t,\tau)=(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T},$ (2)
-where $\mathrm{E}_{t}$ denotes the conditional expectation given $e_{t}$.
-We will see that $\tau^{*}\in\mathcal{T}$ could be an optimal control function for any initial period $t$ and for later specified $\{\lambda_{k}\}_{k=t}^{\infty}$ such that
-$J(\lambda_{t},t,\tau^{*})=\sup_{\tau\in\mathcal{T}}$ $(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right].$ (3)
-2.2 Pareto optimal transfer mechanism with asymmetric information
-If income is private information, the government has to ask individuals to declare their income.
+$v_{1}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathbb{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})]=u((u^{\prime})^{-1}(\lambda_{0}^{-1}))=:\overline{v}_{1}(\lambda_{0})\,\text{(promised utility)}$
+$v_{2}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathbb{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}(-\tau_{t})]\,\text{(expected budget impact)}.$
+The map $\lambda\mapsto\overline{v}_{1}(\lambda)$ is strictly increasing and will serve as the state variable once income becomes private. We also write
+$\overline{v}_{2}(\lambda_{t}):=\mathbb{E}[(1-\beta)\mathbb{E}_{t}[\sum_{s=t}^{\infty}\beta^{s-t}(-\tau^{\ast}(e_{s}|\lambda_{s}))]].$
+Remark 1.
+(Lifetime value functional) Let $\mathcal{T}$ be the set of all admissible control functions.
+Given an initial promise $\lambda_{t}$ and any admissible transfer rule $\tau\in\mathcal{T}$, each individual’s discounted lifetime utility is
+$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
+This function will be used in Section 5 to verify that our mechanism solves the associated Bellman equation.
+2.3 Private information and mechanism design
+If income is private information, the planner has to ask individuals to declare their income.
 In general, a mechanism $\Gamma$ that arranges transfers $\{\tau_{t}\}_{t=0}^{\infty}$ needs to satisfy sequential incentive compatibility defined below.
-For this definition, let $z$ denote the current observable state, which can be a finite vector of state variables, and may include the past history of observables or be in a recursive form.
-The distribution of the tomorrow’s state $z^{\prime}$ is determined by this $z$ and the current level of consumption $x$.
-We write $\pi(x)$ for the value of each individual’s current payoff and $v(z)$ for the present value of the mechanism $\Gamma$ at $z$.
 Definition 1.
-[Sequential incentive compatibility, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be sequentially incentive compatible if for every sate $z$
-$v(z)=(1-\beta)\pi(x)+\beta\mathrm{E}_{(x,z)}[v(z^{\prime})]\geq(1-\beta)\pi(\tilde{x})+\beta\mathrm{E}_{(\tilde{x},z)}[v(z^{\prime})],$
-where $E_{(x,z)}$ means that the conditional expectation given $(x,z)$.
+[Sequential incentive compatibility, Marcet and Marimon, 1992] Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
+$(1-\beta)u(e+\tau(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\geq(1-\beta)u(e+\tau(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\quad\forall e,\tilde{e}\in E.$
 We further evaluate a mechanism according to the following notion.
 Definition 2.
-[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequnetially incentive compatible mechanism.
-In addition to taking these things into account, since it is known to lead to an immisarization outcome, the government does not hope to manipulate the transfer function $\tau^{*}$.
-The government takes an alternative approach.
-It is to vary the utility control variable $\lambda$ in each period.
+[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be a sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequentially incentive compatible mechanism.
+In addition to taking these things into account, since it is known to lead to an immiseration outcome, the planner does not hope to manipulate the transfer function $\tau^{*}$.
+The planner takes an alternative approach.
+It is to vary the promise $\lambda$ in each period.
 This could be an incentive to tell the truth.
-For a given $\lambda$ in a given period, the government promises to ensure that the lifetime utility level $\overline{v}_{1}(\lambda)$, and the utility control variable for the next period is renewed taking into account the current utility control variable and the income reported in the current period.
+For a given $\lambda$ in a given period, the planner promises to ensure that the lifetime utility level $\overline{v}_{1}(\lambda)$, and the promise for the next period is renewed taking into account the current promise and the income reported in the current period.
 If an individual truthfully reports income $e$ in a period $t$, then the lifetime utility from that period should be
-$\overline{v}_{1}(\lambda)=\mathrm{E}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]},$ (4)
-where $\lambda^{\prime}(\cdot|\lambda):E\to\mathbb{R}_{++}$ is the utility control variable for the next period.
-If $\lambda^{\prime}$ satisfies (4), then $\mathrm{E}[\overline{v}_{1}(\lambda^{\prime}(e|\lambda)]=\overline{v}_{1}(\lambda)$.
-Truth telling is then induced if, given the current utility control variable $\lambda$, the renewed utility control variable $\lambda^{\prime}$ satisfies the following incentive constraints.
+$\overline{v}_{1}(\lambda)=\mathbb{E}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]},$ (3)
+where $\lambda^{\prime}(\cdot|\lambda):E\to\mathbb{R}_{++}$ is the promise for the next period.
+If $\lambda^{\prime}$ satisfies (3), then $\mathbb{E}[\overline{v}_{1}(\lambda^{\prime}(e|\lambda)]=\overline{v}_{1}(\lambda)$.
+Truth-telling is then induced if, given the current promise $\lambda$, the renewed promise $\lambda^{\prime}$ satisfies the following incentive constraints.
 $(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))$
 $\geq(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\text{ for }e,\tilde{e}\in E.$ (4)
-Such an utility control variable $\lambda^{\prime}$ that satisfies both (4) and (5) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$.
-By a feasible $\lambda^{\prime}$ we mean that the updated utility control variable keeps the level of expected utility for the next period below the supremum of each individual’s utility function.
+Such a promise $\lambda^{\prime}$ that satisfies both (3) and (4) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$.
+By a feasible $\lambda^{\prime}$ we mean that the updated promise keeps the level of expected utility for the next period below the supremum of each individual’s utility function.
 That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$.
@@
 $\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))\big{)}.$ (5)
-Since $v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)=(1-\beta)(e-\mathrm{E}[e_{t}])$, if the reported income $e\in E$ is above the average income, $\mathrm{E}[e_{t}]$, this will be rewarded with higher utility control variables for the next period.
-Conversely, if the reported income $e\in E$ is below the average income, $\mathrm{E}[e_{t}]$, this will result in lower utility control variables for the next period.
-Note that $u(e+\tau^{*}(e|\lambda))=u((u^{\prime})^{-1}(\lambda^{-1}))=\overline{v}_{1}(\lambda)$.
-From this we see that $\lambda^{\prime}$ defined in (6) satisfies (4):
-$\mathrm{E}[(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))]$
-$=\mathrm{E}[\overline{v}_{1}(\lambda)+\lambda^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))]=\overline{v}_{1}(\lambda).$
-Furthermore, remind that the government’s objective function in problem (1) with $\lambda_{0}=\lambda$:
-$\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)$
-is optimal at true income report $e$. It follows that
-$\lambda\big{[}(1-\beta)$ $u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)$
-$\geq\lambda[(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda)]+v_{2}(\lambda,\tilde{e})-\overline{v}_{2}(\lambda)\quad\text{ for }e,\tilde{e}\in E.$
-Then we have the following inequality and see that (5) holds.
-$\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)$
-$=\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]}$
-$\geq\lambda[(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda)]+v_{2}(\lambda,\tilde{e})-\overline{v}_{2}(\lambda)$
-$=\lambda\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\big{]}\quad\text{ for }e,\tilde{e}\in E.$
-On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (6) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the government must find the closest utility control variable for which there is a certainty equivalent to the promised level of lifetime utility.
+Under the updating rule of $\lambda^{\prime}$ in (5), misreporting never raises the weighted sum $\lambda u(e+\tau^{*}(e|\lambda))+v_{2}(\lambda,e)$, hence truth-telling is optimal.
+On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (5) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the planner must find the closest promise for which there is a certainty equivalent to the promised level of lifetime utility.
 That is, for given $\lambda$, if there is $\hat{e}\in E$ such that $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\hat{e})-\overline{v}_{2}(\lambda))\geq S$, for reported income $e\in E$, define $\lambda^{\prime}$ such that
-$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},\overline{e}_{\lambda}\})])\big{)},$ (7)
-where $\overline{e}_{\lambda}$ is defined such that
-$\overline{e}_{\lambda}=\arg\max_{e^{\prime}\in E_{N}}\big{(}v_{2}(\lambda,e^{\prime})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}$
+$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\bigg{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})-\mathbb{E}_{\tau^{\ast}}\left[v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})\right]\bigg{)},$ (6)
+where $\mathbb{E}_{\tau^{\ast}}$ denotes
+$\mathbb{E}_{\tau^{\ast}}\left[v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})\right]:=(1-\beta)\mathbb{E}\left[-\tau^{\ast}(\min\{e_{t},\overline{e}_{\lambda}\}|\lambda)+\sum_{s=t+1}^{\infty}\beta^{s-t}(-\tau^{\ast}(e_{s}|\lambda_{s}))\right]$
+$\overline{e}_{\lambda}$ is defined such that
+$\overline{e}_{\lambda}=\arg\max_{e^{\prime}\in E}\big{(}v_{2}(\lambda,e^{\prime})-\mathbb{E}_{\tau^{\ast}}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}$
 subject to
-$\overline{v}_{1}(\lambda)+\frac{1}{\lambda\beta}\big{(}v_{2}(\lambda,e^{\prime})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}<S,$
-where $E_{N}\supset E$ is a set with $|E_{N}|=N>M$, $N\in\mathbb{N}$, and for $e\in E_{N}$ $e^{1}\leq e\leq e^{M}$.
-[^2] To avoid getting stuck on a pass where the same utility control variable $\lambda$ goes on forever, if you find $\overline{e}_{\lambda}=e^{1}$ for given $N$, choose $N^{\prime}\in\mathbb{N}$ with $N^{\prime}>N$, and use $E_{N^{\prime}}$ instead of $E_{N}$ so as to obtain $\overline{e}_{\lambda}\geq e^{2}$.
-This adjustment also changes the transfers in the current period $t$.
-However, we see the recursive relation in (7) satisfies the following corresponding incentive constraint.
+$\overline{v}_{1}(\lambda)+\frac{1}{\lambda\beta}\big{(}v_{2}(\lambda,e^{\prime})-\mathbb{E}_{\tau^{\ast}}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}<S.$
+This adjustment also changes the transfers in the current period $t$. However, we see the recursive relation in (6) satisfies the following corresponding incentive constraint.
 $(1-\beta)u(e$ $+\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))$
@@
 This is because the transfer $\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda)$ in (7) is the solution to the problem (1) subject to $\tau(e_{t})\geq(u^{\prime})^{-1}(\lambda^{-1})-\overline{e}_{\lambda}$ for $t\in\mathbb{N}\cup\{0\}$, where $\lambda=\lambda_{0}$.
-In other words, since the transfer to the government is constant for individuals whose income is above the threshold $\overline{e}_{\lambda}$, they are indifferent between reporting and misreporting their true income.
-With respect to the promise-keeping constraint, it holds because the expectation of the left-hand side of the inequality in (8) is greater than or equal to the promised level of lifetime utility $\overline{v}_{1}(\lambda)$:
-$\mathrm{E}\big{[}(1-\beta)u(e+\tau^{*}(\min(e,\overline{e}_{\lambda})|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime})\big{]}\geq\overline{v}_{1}(\lambda).$
+In other words, since the transfer to the planner is constant for individuals whose income is above the threshold $\overline{e}_{\lambda}$, they are indifferent between reporting and misreporting their true income.
+Concerning the promise-keeping constraint, it holds because the expectation of the left-hand side of the inequality in (7) is greater than or equal to the promised level of lifetime utility $\overline{v}_{1}(\lambda)$:
+$\mathbb{E}\big{[}(1-\beta)u(e+\tau^{*}(\min(e,\overline{e}_{\lambda})|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime})\big{]}\geq\overline{v}_{1}(\lambda).$
+3 Risk-shifting mechanism
+This section builds and analyzes the transfer rule $\mathcal{M}$.
+We first show that each individual’s promised utility forms a bounded martingale, and then derive conditions under which the planner’s inter-temporal budget is non-negative.
 Since the definition of $\lambda^{\prime}$ in (6) is an extension of (5), we can use it for the case where $\lambda^{\prime}$ could also be defined in (5). We define a mechanism like this.
 Definition 3.
-An inter-period transfer mechanism $\mathcal{M}$ is represented by a sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ which satisfies (7) and the corresponding transfers, $\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$.
-A sufficient condition for the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (6) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal transfer $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$.
+(Risk-shifting mechanism) A mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6).
+Within each period $t$, consumption is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1})$; schocks affect only $\lambda_{t+1}$.
+A sufficient condition for the sequence of promise $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (5) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal consumption $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$.
+Assumption 1.
+(Moderate risk aversion). For all $\lambda>0$,
+$-\frac{u^{\prime\prime}}{u^{\prime}}((u^{\prime})^{-1}(\lambda^{-1}))<\frac{\beta}{(1-\beta)(e^{M}-\mathbb{E}[e_{t}])}.$ (8)
+Interpretation.
+Arrow-Pratt risk aversion at the contract consumption level is smaller than a scaled income spread; the condition is mild when $\beta$ is close to one.
+Under Assumption 1, feasibility does not bind in the promise-update map, and the recursive rule (5) applies period by period.
+Remark 2.
+We could remove condition Assumption 1 by removing the assumption of bounded utility; dropping it would not violate sequential efficiency but would complicate the martingale proof.
+Boundedness is used to ensure that $\{\overline{v}_{1}(\lambda_{t})\}$ converges.
+Lemma 1.
+Under Assumption 1 the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (5).
+Proof: see Appendix A.
+If the inequality (8) does not hold, there may be a case where a promised-utility state $\lambda_{t}$ in $\mathcal{M}$ is defined in (6) instead of in (5), e.g.
+if the individuals’ utilities have constant absolute risk aversion, and if it is greater than $\beta/((1-\beta)(e^{M}-\mathbb{E}[e_{t}]))$.
+The risk-shifting mechanism $\mathcal{M}$ is sequentially incentive compatible regardless of whether each $\lambda_{t}$ is defined in (5) or in (6).
+Furthermore, if $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (5), like the $\lambda$ mechanism of Marcet and Marimon [1992], $\mathcal{M}$ is sequentially efficient.
+Theorem 1.
+The mechanism $\mathcal{M}$ defined by $(\tau^{\ast},\{\lambda_{t}\})$ is sequentially incentive compatible. Under Assumption 1, it is sequentially efficient.
+Proof: see Appendix A.
+3.1 Individual level: bounded promised utilities
+This subsection shows that the sequence of promised utilities forms a bounded martingale and therefore converges, so immiseration cannot occur.
+Let $(\mathcal{F}_{t})$ be the natural filtration generated by $\{\lambda_{s}\}_{s\leq t}$.
+Because (6) makes $\lambda_{t+1}$ measurable w.r.t.
+the $\sigma$-algebra generated by $\lambda_{t}$ and $e_{t}$, the process $\overline{v}_{1}(\lambda_{t})$ is adapted.
+Theorem 2.
+The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a bounded martingale and converges to an integrable random variable a.s.
+(no immiseration).
+Proof: see Appendix A.
+Because promised utility is a martingale with bounded support, the mechanism is a fair intertemporal lottery: expected lifetime utility is preserved ex-ante, but its realisation drifts neither up nor down on average.
+3.2 Planner’s budget and sustainability
+All claims in this subsection are sufficient (not necessary) conditions based on a second-order expansion.
+We derive sufficient conditions under which the cross-sectional budget is non-negative each period.
+The convergence of individuals’ lifetime utilities to an integrable random variable with finite expectation does not yet ensure the sustainability of the $\mathcal{M}$ mechanism.
+This is because the planner predicts a declining trend in its budget for each individual in the long run, as we see in the following lemma.
+Let $\overline{\tau}_{n}$ denote the average transfer to an $n$-period-old individual (positive means the planner pays).
+Then $\overline{\tau}_{n}$ is:
+$\overline{\tau}_{n}$ $=\mathbb{E}[\tau^{\ast}(e_{n}|\lambda_{n})]=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathbb{E}[e_{n}]\quad\text{ for }n\in\mathbb{N}\cup\{0\}.$
+Lemma 2.
+The average transfer for each cohort $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is an increasing sequence.
+Proof sketch (details in Appendix B).
+For $n=1$, since $\mathbb{E}[\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
+$\overline{\tau}_{1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathbb{E}[e_{1}]$
+$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathbb{E}[e_{0}]=\overline{\tau}_{0}.$
+The proof of Lemma 2 shows that the increasing trend in $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is due to the cost of keeping the expected lifetime utility in each period unchanged from the previous period, although there are risks between periods.
+Let $g_{t}(\lambda_{0})$ be the expected budget balance in period $t$ for the planner for a given initial promise $\lambda_{0}$.
+Since each individual’s probability of being alive is $\alpha\in(0,1)$, and from period one the next generation of continuum of individuals on the $(1-\alpha)$ interval is born in each period, $g_{t}(\lambda_{0})$ is given by
+$g_{t}(\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}.$
+Here, a positive $g_{t}$ denotes a surplus (the planner collects funds).
+For the $\mathcal{M}$ mechanism to be sustainable, the budget balance in each period must be greater than or equal to zero.
+As we see in the following proposition, under a second-order (Taylor) approximation, if incomes have a symmetric distribution, a sufficient condition for the budget balance in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
+Theorem 3.
+In a stationary overlapping-generations society with i.i.d. income risk, if
+  • (i)
+    incomes have a symmetric distribution,
+  • (ii)
+    $u^{\prime\prime}/u^{\prime}$ is a decreasing function,
+  • (iii)
+    $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function and
+  • (iv)
+    initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied,
+then planner’s period-by-period balance $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ under a second-order (Taylor) approximation.
+Proof: Second-order approximation shows decreasing differences. See Appendix C.
+4 Numerical examples
+An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Theorem 3 is given by the second-order polynomial approximation as follows.
+Corollary 1.
+The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Theorem 3 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
+$\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))\geq\sqrt{\frac{(1-\alpha r)^{2}}{\alpha r^{2}(1-\alpha)}\mathrm{Var}[e_{t}]},$ (9)
+where $\lambda_{E}=1/(u^{\prime}(\mathbb{E}[e_{t}]))$ denotes the promise corresponding to mean income report.
+Proof: see Appendix D.
+Figure 1: Utility Function
+{singlespace*}
+Notes: This figure illustrates a utility function (vertical axis) against consumption (horizontal axis).
+The segment $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$ appears in Corollary 1.
+It is the linear approximation of the length between the expected income and the consumption guaranteed in the mechanism M with $\lambda_{0}$.
+The left-hand side of the inequality (9) is a linear approximation of the difference between the expected income and the consumption level in the mechanism $\mathcal{M}$ with the promise $\lambda_{0}$.
+The inequality (9) means that this length is greater than the square root of the weighted income variance.
+Figure 1 shows which part of the graph of a utility function corresponds to $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$.
+The lifetime utility $\overline{v}_{1}(\lambda_{0})$ in the mechanism $\mathcal{M}$ is higher than that in the state of autarky if the risk premium of the state of autarky is higher than the term in the right-hand side of the inequality (9).
+A higher discount factor and higher probability of being alive will relax the sufficient condition of $\lambda_{0}$ for making the mechanism $\mathcal{M}$ sustainable.
+If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (9) may be unnecessarily strict.
+We illustrate the mechanism with a CRRA utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
+  • Utility $u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma}$, $\gamma=-3$
+  • Income support $E=\{0,1,2,3,4,5,6,7,8,9\}:=E_{s}$
+  • Probabilities $P(e)$ (mirror-symmetric, approx. normal) $P_{s}$
+  • Discount factor and survival rate $r=\alpha=0.93$
+  • Initial state $\lambda_{0}=50\in[\lambda_{CE},133]$
+The values follow Corollary 1: the lower bound $\lambda_{CE}=1/u^{\prime}\bigl{(}E[u(e_{t})]\bigr{)}=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
+Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals of the same generation and corresponding transfer balances for $100$ periods for the above case.
+The initial promise is set $\lambda_{0}=50$.
+Consistent with Theorem 2, the utility path stays bounded while transfers $\{-\overline{\tau}_{n}\}_{n=0}^{100}$ fall, echoing Lemma 2.
+Figure 2: Sample mean of lifetime utilities and transfer balances
+Figure 3 shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods.
+In this society, each individual will be alive in the next period with probability $\alpha=0.93$, and newly born individuals enter the society, holding the total population $100$.
+As in Figure 2, each individual’s utility is $u(c)=(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial promise $\lambda_{0}=50$ and the discount factor $r=0.93$.
+The income set $E$ is $E_{s}$.
+The random numbers are generated with the same symmetric distribution $P_{s}$.
+The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and supports Theorem 3.
+Figure 3: Sample means of budget balance $\{g_{t}(\lambda_{0})\}$
+Remark 3.
+Classic hidden-income contracts [e.g.
+Green, 1987, Thomas and Worrall, 1990] spread shocks within each period: truth telling is rewarded by a higher spot transfer, so lifetime utilities diverge over time and immiseration follows.
+Under our mechanism, by contrast, the realized income $e_{t}$ only determines next period’s promise $\lambda_{t+1}$; consumption in period $t$ is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda^{-1}_{t})$ with probability $1$.
+The shock is therefore shifted one step forward—a transfer of risk to the future.
+This intertemporal insurance does make the average transfer profile $\{-\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ fall with age, threatening the planner’s long-run budget.
+Theorem 3 shows the problem can be neutralized by choosing an initial promise $\lambda_{0}$ that gives early cohorts negative net transfers, so later positive transfers are prefunded.
+Finally, note a key difference from Marcet and Marimon [1992]: in their growth model, investment returns are themselves the incentive, and the contract is bilateral and non-competitive, which can depress the manager’s utility below autarky.
+Our pure-exchange setting has no such external scale effect, so every cohort attains a higher lifetime utility than autarky.
+5 Dynamic-programming verification
+This section verifies that the transfer rule $\tau^{\ast}$ solves the infinite-horizon Bellman equation and that $V$ is the corresponding value function (under condition (8) of Lemma 1).
+The Bellman equation (DP) is given by
+$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)]\},$ (10)
+where $\Lambda(\lambda_{t})$ is a set of measurable controls satisfying incentive and promise-keeping, and $\lambda_{t+1}$ is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{R}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (6).
+Our candidate solution is $V$ in (11) with control $\tau^{\ast}$.
+$V(\lambda_{t},t)=\mathbb{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]},\quad\text{for }\,\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}.$ (11)
+We observe the following proposition.
+Proposition 1.
+With Assumption 1, $V$ in (11) solves the infinite-horizon Bellman equation, and $\tau^{*}$ maximizes the lifetime value $J$ in (2).
+Proof sketch.
+[…]
+(complete proof in Appendix E).
+6 Conclusion
+This paper shows that the grim ”immiseration“ prediction of Green [1987] and Thomas and Worrall [1990] is not inevitable.
+A simple risk shifting rule that postpones risk can deliver near full within-period insurance while keeping promised utilities bounded.
+Provided the initial promise satisfies a transparent inequality, inter-generational cooperation sustains the mechanism, and every cohort attains higher lifetime utility than under autarky, with no within-cohort inequality.
+Appendix A Proof of propositions 1 and 2
+Proof of Lemma 1.
 Write $F(\lambda)=\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))$ for $\lambda>0$.
-Lemma 1.
-The sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (6) if
-$-\frac{u^{\prime\prime}((u^{\prime})^{-1}(\lambda^{-1}))}{u^{\prime}((u^{\prime})^{-1}(\lambda^{-1}))}<\frac{\beta}{(1-\beta)(e^{M}-\mathrm{E}[e_{t}])}\quad\text{for}\quad\lambda>0.$ (9)
-Proof.
 It suffices to show the proposition that $F(\lambda)<S$ for all $\lambda>0$.
@@
 ∎
-If condition (9) does not hold, there may be a case where an utility control variable $\lambda_{t}$ in $\mathcal{M}$ is defined in (7) instead of in (6), e.g.
-if the individuals’ utilities have constant absolute risk aversion, and if it is greater than $\beta/((1-\beta)(e^{M}-\mathrm{E}[e_{t}]))$.
-Since $F$ decreases in $\lambda$ and $\lim_{\lambda\to 0}F(\lambda)=\infty$, there must be a threshold $\overline{\lambda}$ below which $F(\lambda)\geq S$.
-The inter-period transfer mechanism $\mathcal{M}$ is sequentially incentive compatible regardless of whether each $\lambda_{t}$ is defined in (6) or in (7).
-Furthermore, if $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (6), like the $\lambda$ mechanism of Marcet and Marimon [1992], $\mathcal{M}$ is sequentially efficient.
-Proposition 1.
-$\mathcal{M}$ is sequentially incentive compatible. Furthermore, if condition (9) holds, then $\mathcal{M}$ is sequentially efficient.
-Proof.
+Proof of Theorem 1.
 With respect to $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ defined throughout in (5), the propositional statement is shown by Marcet and Marimon [1992]. For completeness, we report the entire proof.
@@
 $\lambda_{t}\big{[}$ $(1-\beta)u(e+\tau^{*}(e|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e|\lambda_{t}))\big{]}$
-=(6)${=}\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e)-\overline{v}_{2}(\lambda_{t})$ )
+[…]
 $\geq\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},\tilde{e})-\overline{v}_{2}(\lambda_{t})$
@@
 $\lambda_{t}\big{[}(1-\beta)$ $u(e+\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t}^{\prime}(e|\lambda_{t}))\big{]}$
-=(7)${=}\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}$
-$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{e,\overline{e}_{\lambda_{t}}\})-\mathrm{E}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
+[…]
+$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{e,\overline{e}_{\lambda_{t}}\})-\mathbb{E}_{\tau^{*}}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
 $\geq\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}$
-$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{\tilde{e},\overline{e}_{\lambda_{t}}\})-\mathrm{E}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
+$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{\tilde{e},\overline{e}_{\lambda_{t}}\})-\mathbb{E}_{\tau^{*}}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
 $=\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t}^{\prime}(\tilde{e}|\lambda_{t}))\big{]}\text{ for }e,\tilde{e}\in E.$
-The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
+The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) (planner’s static problem at given $\lambda$) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
 We proceed to show that if condition (8) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms.
 By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible.
-By Lemma 1 since condition (9) holds, the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (6).
+By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (5).
 Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period.
@@
 Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$.
-However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under condition (9).
+However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under the condition (8).
 ∎
-Proposition 1 shows sequential efficiency of the mechanism $\mathcal{M}$ only for the case where condition (9) holds.
-However, we could relax this restriction by removing the assumption that the utility function is bounded from above.
-This assumption provides a sufficient condition to ensure Proposition 2, but it is not a necessary condition.
-3 The inter-period transfer mechanism for each individual
-The government’s approach of using the mechanism $\mathcal{M}$ successfully avoids an immiserization outcome.
-That is, the promised lifetime utility of each individual in $\mathcal{M}$ converges almost everywhere to a random variable with a finite expectation.
-To see this, we reinterpret $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}}$ as a sequence of random variables on $(\Omega,\mathcal{F},P)$.
-This is done recursively as follows.
-First we interpret $\overline{v}_{1}(\lambda_{1})$ as a random variable $\overline{v}_{1}(\lambda_{1}(e_{0}(\cdot)|\lambda_{0}))$ and let $\mathcal{F}_{1}\subset\mathcal{F}$ be the smallest $\sigma$-algebra induced by $\overline{v}_{1}(\lambda_{1})$.
-That is, $\mathcal{F}_{1}=\{\overline{v}_{1}(\lambda_{1})^{-1}(B)\ |\ B\in\mathcal{B}(\mathbb{R})\}$, where $\overline{v}_{1}(\lambda_{1})^{-1}(B)=\{\omega\in\Omega\ |\ \overline{v}_{1}(\lambda_{1}(e_{0}(\omega)|\lambda_{0}))\in B\}$.
-We see that $\overline{v}_{1}(\lambda_{1})$ is $\mathcal{F}_{1}$-measurable, $\overline{v}_{1}(\lambda_{1}):(\Omega,\mathcal{F}_{1})\to(\mathbb{R},\mathcal{B}(\mathbb{R})).$ Second, we consider $\overline{v}_{1}(\lambda_{n})$ as a random variable $\overline{v}_{1}(\lambda_{n}(e_{n-1}(\cdot)|\lambda_{n-1}))$ and define $\mathcal{F}_{n}\subset\mathcal{F}$ as the smallest $\sigma$-algebra induced by the product of the elements in $\{\overline{v}_{1}(\lambda_{i})\}^{n}_{i=1}$ such that
-$\mathcal{F}_{n}=\{(\overline{v}_{1}(\lambda_{1})\times\dots\times\overline{v}_{1}(\lambda_{n}))^{-1}(B)\ |\ B\in\mathcal{B}(\mathbb{R}^{n})\},$
-where
-$(\overline{v}_{1}(\lambda_{1})\times\dots\times$ $\overline{v}_{1}(\lambda_{n}))^{-1}(B)$
-=$=\Big{\{}\omega\in\Omega\ \Big{|}\ \big{(}\overline{v}_{1}(\lambda_{1}(e_{0}(\omega)|\lambda_{0})),\dots,\overline{v}_{1}(\lambda_{n}(e_{n-1}(\omega)|\lambda_{n-1}))\big{)}\in B\Big{\}}.$
-We see that $\overline{v}_{1}(\lambda_{n})$ is $\mathcal{F}_{n}$-measurable, $\overline{v}_{1}(\lambda_{n}):(\Omega,\mathcal{F}_{n})\to(\mathbb{R},\mathcal{B}(\mathbb{R})).$
-Proposition 2.
-The sequence of promised lifetime utilities of individuals $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ converges to an integrable random variable.
-Proof.
+Proof of Theorem 2.
 We first show that $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale. Since the sequence $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}\in\mathcal{M}$ satisfies (6), we have
-$\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})(\omega)]=\overline{v}_{1}(\lambda_{n-1})(\omega)$
-for given realised $\overline{v}_{1}(\lambda_{n-1})(\omega)$. We thus have the following equalities: For all $B\in\mathcal{B}(\mathbb{R})$ for all $n\in\mathbb{N}$
-[…]
+$\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})(\omega)]=\overline{v}_{1}(\lambda_{n-1})(\omega)$
+for given realized $\overline{v}_{1}(\lambda_{n-1})(\omega)$. We thus have the following equalities: For all $B\in\mathcal{B}(\mathbb{R})$ for all $n\in\mathbb{N}$
+$\int_{B}\mathbb{E}\big{[}\overline{v}_{1}(\lambda_{n})\big{|}\overline{v}_{1}(\lambda_{n-1})(\omega)\big{]}dP_{\overline{v}_{1}(\lambda_{n-1})}$ $=\int_{B}\overline{v}_{1}(\lambda_{n-1})(\omega)dP_{\overline{v}_{1}(\lambda_{n-1})},$
 and
-[…]
-[…]
-From the last equality we have for $n\in\mathbb{N}$ $\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e.
+$\int_{B}\mathbb{E}\big{[}\overline{v}_{1}(\lambda_{n})\big{|}\overline{v}_{1}(\lambda_{n-1})(\omega)\big{]}dP_{\overline{v}_{1}(\lambda_{n-1})}$ $=\int_{\{\overline{v}_{1}(\lambda_{n-1})(\omega)\in B\}}\overline{v}_{1}(\lambda_{n})dP$
+$=\int_{\{\overline{v}_{1}(\lambda_{n-1})(\omega)\in B\}}\overline{v}_{1}(\lambda_{n-1})dP.$
+From the last equality we have for $n\in\mathbb{N}$ $\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e.
 $[P]$, and $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale as claimed.
@@
 So we have $\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$.
-Since $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale and $\sup_{n}\mathrm{E}[\overline{v}_{1}(\lambda_{n})^{+}]\leq\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$, from the submartingale convergence theorem, there is an integrable random variable $v^{\infty}$ such that $\overline{v}_{1}(\lambda_{n})\to v^{\infty}$ almost everywhere.
+Since $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale and $\sup_{n}\mathbb{E}[\overline{v}_{1}(\lambda_{n})^{+}]\leq\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$, from the submartingale convergence theorem, there is an integrable random variable $v^{\infty}$ such that $\overline{v}_{1}(\lambda_{n})\to v^{\infty}$ almost everywhere.
 ∎
-As the proof of Proposition 2 shows, mechanism $\mathcal{M}$ functions like a fair-value gamble on lifetime utilities across periods.
-The lifetime utility in each period is a random variable whose sequence forms a martingale.
-Consequently, the mechanism is fair in the sense that each individual’s expected lifetime utility remains unchanged from the time of the bet.
-4 The inter-period transfer mechanism for the government
-However, the convergence of individuals’ lifetime utilities to an integrable random variable with finite expectation does not yet ensure the sustainability of the $\mathcal{M}$ mechanism.
-This is because the government predicts a declining trend in its budget for each individual in the long run, as we see in the following lemma.
-Let $\overline{\tau}_{n}$ be the average transfer to those who were born $n$ periods ago.
-Then $\overline{\tau}_{n}$ is given by as follows.
-$\overline{\tau}_{n}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathrm{E}[e_{n}]\quad\text{ for }n\in\mathbb{N}\cup\{0\}.$
-Lemma 2.
-The average transfer for each age of individuals $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is an increasing sequence.
-Proof.
-We first consider the case $n=1$. For $n=1$, since $\mathrm{E}[\overline{v}_{1}(\lambda_{1}(e_{1}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
-$\overline{\tau}_{1}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathrm{E}[e_{1}]$
-$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathrm{E}[e_{0}]=\overline{\tau}_{0}.$
-For $n>1$ since $u^{-1}$ is a convex function, from Jensen’s inequality we see the result.
-$\overline{\tau}_{n-1}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n-1}))]-\mathrm{E}[e_{n-1}]$
-$=\mathrm{E}[u^{-1}(\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})])]-\mathrm{E}[e_{n-1}]$
-$\leq\mathrm{E}[\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))|\overline{v}_{1}(\lambda_{n-1})]]-\mathrm{E}[e_{n-1}]$
-$=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathrm{E}[e_{n}]$
+Appendix B Proof of Lemma 2
+Proof of Lemma 2.
+We first consider the case $n=1$. For $n=1$, since $\mathbb{E}[\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
+$\overline{\tau}_{1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathbb{E}[e_{1}]$
+$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathbb{E}[e_{0}]=\overline{\tau}_{0}.$
+For $n>1$, since $u^{-1}$ is a convex function, from Jensen’s inequality we see the result.
+$\overline{\tau}_{n-1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n-1}))]-\mathbb{E}[e_{n-1}]$
+$=\mathbb{E}[u^{-1}(\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})])]-\mathbb{E}[e_{n-1}]$
+$\leq\mathbb{E}[\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))|\overline{v}_{1}(\lambda_{n-1})]]-\mathbb{E}[e_{n-1}]$
+$=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathbb{E}[e_{n}]$
 $=\overline{\tau}_{n}\quad\text{ for }n=2,3,\cdots.$
-The proof of Lemma 2 shows that the increasing trend in $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is due to the cost of keeping the expected lifetime utility in each period unchanged from the previous period, although there are risks between periods.
-Let $g(t,\lambda_{0})$ be the expected balance of payments in period $t$ for the government for given initial utility control variable $\lambda_{0}$.
-Since each individual’s probability of being alive is $\alpha\in(0,1)$, and from period one the next generation of contimuum of individuals on the $(1-\alpha)$ interval is born in each period, $g(t,\lambda_{0})$ is given by
-$g(t,\lambda_{0})$ $=\alpha^{t}(-\overline{\tau}_{t})+\alpha^{t-1}(1-\alpha)(-\overline{\tau}_{t-1})+\cdots+\alpha(1-\alpha)(-\overline{\tau}_{1})+(1-\alpha)(-\overline{\tau}_{0})$
-$=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}.$
-For the $\mathcal{M}$ mechanism to be sustainable, the balance of payments in each period must be greater than or equal to be zero.
-As we see in the following proposition, with the approximation of the second order Taylor polynomial, if incomes have a symmetric distribution, a sufficient condition for the balance of payments in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
-Proposition 3.
-Assume that (i) incomes have a symmetric distribution, (ii) $-u^{\prime\prime}/u^{\prime}$ is a decreasing function, (iii) $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function or equivalently,
-$\frac{\left(-\frac{u^{\prime\prime}}{u^{\prime}}\right)^{{}^{\prime\prime}}({\scriptstyle u^{-1}(y)})}{-\left(-\frac{u^{\prime\prime}}{u^{\prime}}\right)^{\prime}({\scriptstyle u^{-1}(y)})}\leq\frac{(u^{-1})^{\prime\prime}(y)}{(u^{-1})^{\prime}(y)},$
-and (iv) $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$.
-Then $g(t,\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ with the approximation of the second order Taylor polynomial.
-The proof of Proposition 3 uses the following lemma.
-Lemma 3.
-Under the assumptions (i), (ii), (iii) in Proposition 3, the second order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
+Appendix C Proof of Theorem 3
+The proof of Theorem 3 uses the following lemma.
+Lemma C.1.
+Under the assumptions (i), (ii), (iii) in Theorem 3, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
 Proof.
 First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (5).
-Using the third order Taylor polynomial, for a given utility control variable $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
-$u^{-1}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=$ $u^{-1}(\overline{v}_{1}(\lambda_{n-1}))+\frac{1}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\Delta_{n-1}-\frac{1}{2!}\frac{u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{3}}\Delta_{n-1}^{2}$
+Using the third-order Taylor polynomial, for a given promise $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
+$u^{-1}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))-u^{-1}(\overline{v}_{1}(\lambda_{n-1}))=$ $\frac{1}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\Delta_{n-1}-\frac{1}{2!}\frac{u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{3}}\Delta_{n-1}^{2}$
 $+\frac{1}{3!}\left(-\frac{u^{\prime\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{4}}+\frac{3u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{2}}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{5}}\right)\Delta_{n-1}^{3}$
-$+h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))\Delta_{n-1}^{3},$ (10)
-where $\Delta_{n-1}=(1-\beta)(e-\mathrm{E}[e_{n}])/(\beta\lambda_{n-1})$ and $h_{3}:\mathbb{R}\to\mathbb{R}$ is a function such that $\lim_{\Delta_{n-1}\to 0}h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=0$.
-Taking the expectation of (10), $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is written as
-$\overline{\tau}_{n}-\overline{\tau}_{n-1}$ $=\mathrm{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}[e_{n}]\right]+o(\Delta_{n-1}^{3}).$
-Note that since incomes follow a symmetric distribution, the third term on the right-hand side of the equation (10) will cancel out when the expectation is taken.
+$+h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))\Delta_{n-1}^{3},$ (C.1)
+where $\Delta_{n-1}=(1-\beta)(e-\mathbb{E}[e_{n}])/(\beta\lambda_{n-1})$ and $h_{3}:\mathbb{R}\to\mathbb{R}$ is a function such that $\lim_{\Delta_{n-1}\to 0}h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=0$.
+Taking the expectation of (C.1), $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is written as
+$\overline{\tau}_{n}-\overline{\tau}_{n-1}$ $=\mathbb{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}[e_{n}]\right]+o(\Delta_{n-1}^{3}).$
+Note that since incomes follow a symmetric distribution, the third term on the right-hand side of the equation (C.1) will cancel out when the expectation is taken.
 Since $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function, from Jensen’s inequality we have
-$\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\left(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))\right)\right]$ $=\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}\big{(}\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})]\big{)}\big{)}\right]$
-$\geq\mathrm{E}\left[\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n}))\big{)}\bigg{|}\overline{v}_{1}(\lambda_{n-1})\right]\right]$
-$=\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
+$\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\left(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))\right)\right]$ $=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}\big{(}\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})]\big{)}\big{)}\right]$
+$\geq\mathbb{E}\left[\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n}))\big{)}\bigg{|}\overline{v}_{1}(\lambda_{n-1})\right]\right]$
+$=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
 It follows that the third order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence.
-We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (7). We change $\Delta_{n-1}$ in (10) to
-$\overline{\Delta}_{n-1}=\frac{(1-\beta)\big{(}\min\{e,\overline{e}_{\lambda_{n-1}}\}-\mathrm{E}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\big{)}}{\beta\lambda_{n-1}}.$
-By the asymmetric distribution of conditional income $e\leq\overline{e}_{\lambda_{n-1}}$, the third term on the right-hand side of (10) remains and $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is given by
-$\overline{\tau}_{n}-\overline{\tau}_{n-1}=\mathrm{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\right]+o(\overline{\Delta}_{n-1}^{2}).$
-By a similar argument, the second order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
-We prove Proposition 3 below.
-Proof.
-From Lemma 3, with the second order polynomial approximation we have
+We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (6). We change $\Delta_{n-1}$ in (C.1) to
+$\overline{\Delta}_{n-1}=\frac{(1-\beta)\big{(}\min\{e,\overline{e}_{\lambda_{n-1}}\}-\mathbb{E}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\big{)}}{\beta\lambda_{n-1}}.$
+By the asymmetric distribution of conditional income $e\leq\overline{e}_{\lambda_{n-1}}$, the third term on the right-hand side of (C.1) remains and $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is given by
+$\overline{\tau}_{n}-\overline{\tau}_{n-1}=\mathbb{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\right]+o(\overline{\Delta}_{n-1}^{2}).$
+By a similar argument, the second-order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
+We prove Theorem 3 below.
+Proof of Theorem 3.
+From Lemma C.1, with the second-order polynomial approximation, we have
 $\overline{\tau}_{n}-\overline{\tau}_{n-1}\leq\overline{\tau}_{1}-\overline{\tau}_{0}\quad\text{for }n\in\mathbb{N}.$
-Hence the following inequality holds.
-$g(t,\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
+Hence, the following inequality holds.
+$g_{t}(\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
 $\geq\alpha\cdot\frac{1-\alpha^{t}}{1-\alpha}(\overline{\tau}_{0}-\overline{\tau}_{1})-\overline{\tau}_{0}.$
 The last term is positive if
-$-\overline{\tau}_{0}\geq\frac{\alpha(1-\alpha^{t})}{1-\alpha}(\overline{\tau}_{1}-\overline{\tau}_{0})$
-and we see a sufficient condition for $g(t,\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ is $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
-An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Proposition 3 is given by the third order polynomial approximation as follows.
-Corollary 1.
-The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Proposition 3 holds with the second order polynomial approximation if the initial utility control variable $\lambda_{0}$ satisfies
-$\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))\geq\sqrt{\frac{(1-\alpha r)^{2}}{\alpha r^{2}(1-\alpha)}\mathrm{Var}[e_{t}]},$ (11)
-where $\lambda_{E}=1/(u^{\prime}(u(\mathrm{E}[e_{t}])))$.
-Proof.
+$\overline{\tau}_{0}\geq\frac{\alpha(1-\alpha^{t})}{1-\alpha}(\overline{\tau}_{1}-\overline{\tau}_{0})$
+and we see a sufficient condition for $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ is $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
+Appendix D Proof of Corollary 1
+Proof of Corollary 1.
 Let us first examine the case where $\lambda_{1}$ can be defined in (5).
-Applying Tayler theorem to $u^{-1}(\overline{v}_{1}(\lambda_{1}))$, $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)$ is expressed as
-$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{0}]+o(\Delta_{0}^{3}).$ (12)
-On the other hand, applying Tayler theorem to $u^{-1}(\overline{v}_{1}(\lambda_{E}))$, we have
+Applying Taylor theorem to $u^{-1}(\overline{v}_{1}(\lambda_{1}))$, $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)$ is expressed as
+$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{0}]+o(\Delta_{0}^{3}).$ (D.1)
+On the other hand, applying Taylor to $u^{-1}(\overline{v}_{1}(\lambda_{E}))$, we have
 $-\overline{\tau}_{0}$ $=-u^{-1}(\overline{v}_{1}(\lambda_{0}))+u^{-1}(\overline{v}_{1}(\lambda_{E}))$
-$=\frac{(u^{-1})^{\prime}(\overline{v}_{1}(\lambda_{0}))}{1!}\Delta_{E}+\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\Delta_{E}^{2}+o(\Delta_{E}^{2}),$ (13)
+$=\frac{(u^{-1})^{\prime}(\overline{v}_{1}(\lambda_{0}))}{1!}\Delta_{E}+\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\Delta_{E}^{2}+o(\Delta_{E}^{2}),$ (D.2)
 where $\Delta_{E}=\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0})$.
-Since the first term on the right-hand side of (13) is positive, it is sufficient to show the second term on the right-hand side of (13) is equal to or greater than the first term on the righ-hand side of (12).
+Since the first term on the right-hand side of (D.2) is positive, it is sufficient to show the second term on the right-hand side of (D.2) is equal to or greater than the first term on the right-hand side of (D.1).
 However, from the corollary assumption (9), we have
 $\Delta_{E}^{2}\geq\frac{(1-\beta)^{2}}{\lambda_{0}^{2}\beta^{2}}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{t}],$
-which is the desired conclusion that the condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ holds with the second order polynomial approximation.
+which is the desired conclusion that the condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ holds with the second-order polynomial approximation.
 For the case where $\lambda_{1}$ must be defined in (6), we have
@@
 Since $\mathrm{Var}[\min\{e_{0},\overline{e}_{\lambda_{0}}\}]\leq\mathrm{Var}[e_{0}]$, (9) is sufficient to be $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
-Figure 1: Slope of Utility Function
-{singlespace*}
-Notes: This figure shows a graph of the slope of a utility function.
-The horizontal axis is consumption, and the vertical axis is utility.
-The curve is the utility function.
-The length $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$ is a linear approximation of the length between the expected income and the consumption guaranteed in the mechanism M with $\lambda_{0}$.
-The left-hand side of the inequality (11) is a linear approximation of the difference between the consumption level that ensures the expected income and the consumption level in the mechanism $\mathcal{M}$ with the utility control variable $\lambda_{0}$.
-The inequality (11) means that this length is greater than the square root of the weighted income variance.
-Figure 1 shows which part of the graph of a utility function corresponds to $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$.
-The lifetime utility $\overline{v}_{1}(\lambda_{0})$ in the mechanism $\mathcal{M}$ could be set higher than that in the state of autarky if the risk premium of the state of autarky is higher than the term in the right-hand side of the inequality (11).
-A higher discount factor and higher probability of being alive will relax the sufficient condition of $\lambda_{0}$.
-If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (11) may be unnecessarily strict.
-5 Numerical examples
-According to Corollary 1, in the following case, the mechanism $\mathcal{M}$ is sustainable and the lifetime utilities of all generations are higher than that in the state of autarky.
-Suppose that the utility function is given by
-$u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma},$
-where $\gamma=-3$.
-The income set $E$ is $E=\{0,1,2,3,4,5,6,7,8,9\}$.
-We will refer to this income set as $E_{s}$.
-The income shocks follow the distribution $P(e_{t}=0)=Pr(e_{t}=9)=0.025$, $P(e_{t}=1)=P(e_{t}=8)=0.045781$, $P(e_{t}=2)=P(e_{t}=7)=0.092762$, $P(e_{t}=3)=P(e_{t}=6)=0.148524$, $P(e_{t}=4)=Pr(e_{t}=5)=0.187933$.
-We will refer to this distribution as $P_{s}$ Suppose that the discount factor and the probability of being alive are $r=\alpha=0.93$.
-Then Proposition 3 assumptions are satisfied for $\lambda_{0}\in[\lambda_{CE},133]$, where $\lambda_{CE}=49.608=1/u^{\prime}(E[u(e_{t})])$.
-Note that $\lambda_{E}=166.375$, the risk premium is $1.826$, and $\sqrt{(1-\alpha r)^{2}/(\alpha r^{2}(1-\alpha))\mathrm{Var}[e_{t}]}$ is $1.147$.
-Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals and corresponding transfer balances for $100$ periods for the above case.
-The initial utility control variable is set $\lambda_{0}=50$.
-As Proposition 2 implies, the series of sample means of lifetime utilities does not approach negative infinity, which has been called immiserization.
-However, as Lemma 2 shows, the corresponding series of sample means of transfer balances $\{-\tau_{n}\}_{n=0}^{100}$ is decreasing.
-Figure 2: A series of sample means of lifetime utility and corresponding transfer balance
-{singlespace*}
-Notes: This figure shows a series of sample means of lifetime utilities and corresponding transfer balances for $100$ periods, where $1000$ individuals are assumed to have the utility $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$.
-The parameters are set so that the initial utility control variable $\lambda_{0}=50$, the discount factor $r=0.93$, the probability of being alive $\alpha=0.93$.
-The income set $E$ is $E_{s}$.
-The random numbers are generated using the discrete probability distribution $P_{s}$.
-Figure 3 shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods.
-In this society, each individual will be alive in the next period with probalibility $\alpha=0.93$, and newly born individuals enter the society, holding the total population $100$.
-As in Figure 2, each individual’s utility is $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial utility control variable $\lambda_{0}=50$ and the discount factor $r=0.93$.
-The income set $E$ is $E_{s}$.
-The random numbers are generated with the same symmetric distribution $P_{s}$.
-The series of sample means of transfer balances $\{g(n,\lambda_{0})\}_{n=0}^{1000}$ remains non-negative, and supports Proposition 3.
-Figure 3: A series of sample means of balance of payments $\{g(t,\lambda_{0})\}$
-{singlespace*}
-Notes: This figure shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods.
-As in Figure 2, each individual’s utility is $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial utility control variable $\lambda_{0}=50$, the discount factor $r=0.93$, the probability of being alive $\alpha=0.93$.
-The income set $E$ is $E_{s}$.
-The random numbers are generated using the discrete probability distribution $P_{s}$.
-Immiserization observed in previous studies [e.g.
-Green, 1987, Thomas and Worrall, 1990] is due to the approach to allocate risks within each period across the polulation of individuals.
-To ensure truthful reporting, only individuals with the highest incomes keep their lifetime utility level in the next period, and lifetime utilities of other individulas in the next period are discounted according to reported incomes.
-As a result, the average contract value of individuals becomes a decreasing series.
-However, the mechanism $\mathcal{M}$ simply shifts risks to future periods.
-A similar limitation is that, the proposed mechanism leads to a dicreasing series of the average transfer balances for each generation, which could make the government’s balance of payments negative infinity.
-However, this could be solved by setting the initial utility control variable appropriately, so that average transfers are negative for the first few periods from birth.
-If this sense of intergenerational cooperation is properly organised, efficient income allocations could be sustainable even under asymmetric information about income shocks.
-The property of the mechanism $\mathcal{M}$ contrasts with the numerical results of Marcet and Marimon [1992], where the lifetime utility of agent 1 (manager) is higher in the state of autarky than in the state of the incentive contract with agent 2 (investor).
-This is probably because, first, Marcet and Marimon [1992] is a growth model with the assumption that capital and the manager’s consumption, increases when the manager chooses the optimal level of investment.
-Second, since the transfers in the numerical example for incentive contracts in Marcet and Marimon [1992] are not competitive, the external financing contracts end up reducing agent 1’s allocation.
-6 Checking the mecanism optimality for each generation
-Lastly, we express the mechanism $\mathcal{M}$ using a Bellman equation.
-We focus on the case that (9) in Lemma 1.
-First we guess the value function $V:\mathbb{R}_{++}\times\mathbb{N}\cup\{0\}\to\mathbb{R}\cup\{-\infty\}$ is given by for $\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$
-$V(\lambda_{t},t)=\mathrm{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}.$ (14)
-Second, we verify whether this function satisfies the Bellman equation:
-$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1),$ (15)
-where $\Lambda:\mathbb{R}_{++}\times\mathbb{N}\cup\{0\}\multimap\mathbb{U}$ is an available control correspondence defined such that
-$\Lambda(\lambda_{t},t)=\arg\max_{\tau}\left[\lambda_{t}u(e_{t}+\tau)-\tau\ \text{ subject to }(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k}\tau(e_{k})\right]\leq\text{Const.}\right]$. ]
-$\mathbb{U}$ is the set of control parameters, and $\lambda_{t+1}$ in (15) is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{U}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (7).
-We observe the following proposition.
-Proposition 4.
-If the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ satisfies (9), then $V$ fullfills the Bellman equation and it is the value function of the dynamic optmisation problem, where the control function $\tau^{*}$ maximises $J$ in (2).
-A sufficient condition for $V$ to be the value function of the Bellman equation is given by Wiszniewska-Matyszkiel [2011] which is to satisfy the Bellman equation and the following terminal condition.
-Terminal condition [Wiszniewska-Matyszkiel, 2011]
-  • (i)
-    For every $\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$
-    $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}\leq 0$
-    and
-  • (ii)
-    for every $\lambda_{t}\in\{\lambda\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$, if $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}<0$, then
-    $J(\lambda_{t},t,\hat{\tau})=-\infty$
-    for every $\hat{\tau}:\mathbb{R}_{++}\to\mathbb{R}_{+}$ such that $\lambda_{t}$ in $\mathcal{M}$ corresponding to $\hat{\tau}$ and $t$.
-Proof.
-We observe first that if (9) holds, then $V$ defined in (14) satisfies the Bellman equation in (15).
-Indeed, since by definition of $\tau^{*}$ we have $u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))=\overline{v}_{1}(\lambda_{t})$,
-$V(\lambda_{t},t)$ $=\mathrm{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}$
-$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathrm{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1})|\overline{v}_{1}(\lambda_{t})\big{]}$
-$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1).$ (16)
-The last equation follows from the fact that $\lambda_{t+1}$ is realized given $e_{t}$ and seeing $\mathrm{E}_{t+1}$ as a function of $e_{t+1}$. But $\tau^{*}$ is the solution to the problem
-$\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}))+\beta\mathrm{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1})|\overline{v}_{1}(\lambda_{t})\big{]}$ (17)
-under the binding constraint
-$(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k}\tau(e_{k})\right]=\mathrm{E}[\tau(e_{t})]=\text{Const.}$. (18)
-Indeed, for a given $\lambda_{t}$ the objective function in (17) multiplied by $\lambda_{t}$ is written as
-$\lambda_{t}\big{[}(1-\beta)$ $u(e_{t}+\tau(e_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
-$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\mathrm{E}[v_{2}(\lambda_{t},e_{t})]$
-$=(1-\beta)\big{[}\lambda_{t}u(e_{t}+\tau(e_{t}))-\tau(e_{t})\big{]}+\beta\lambda_{t}\overline{v}_{1}(\lambda_{t})+(1-\beta)\mathrm{E}[\tau(e_{t})],$
-and we see that the last terms are maximized by $\tau^{*}$ under the constraint (18). It follows that (16) is written as
-$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1),$
+Appendix E Proof of Proposition 1
+Proof of Proposition 1.
+We observe first that if (8) holds, then $V$ defined in (11) satisfies DP. Indeed, since by definition of $\tau^{*}$ we have $u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))=\overline{v}_{1}(\lambda_{t})$,
+$V(\lambda_{t},t)$ $=\mathbb{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}$
+$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
+$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}\big{[}\mathbb{E}_{t+1}\big{[}\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}\big{]}$
+$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)].$ (E.1)
+But $\tau^{*}$ is the solution to the problem
+$\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)].$ (E.2)
+Indeed, for given information in period $t$, the integrand of the objective function in (E.2) multiplied by $\lambda_{t}$ is written as
+$\lambda_{t}\big{[}(1-\beta)$ $u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
+$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\overline{v}_{2}(\lambda_{t})$
+and we see that the last terms are maximized by $\tau^{*}$. It follows that (E.1) is written as
+$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)],$
 and we see the desired result.
-The task is now to check the terminal condition. Since $u$ is bounded from above, we have
-$\lim_{t\to\infty}$ $V(\lambda_{t},t)\beta^{t}=0.$
-Thus, $V$ satisfies Terminal condition (i) and from Theorem 1 of Wiszniewska-Matyszkiel [2011], we see $V$ defined in (14) is the value function of the Bellman equation in (15), and the control function $\tau^{*}$ maximises $J$ in (2).
+The task is now to check the terminal condition.
+Since $u$ is bounded from above, $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}=0$.
+Hence, the sufficient terminal condition of Wiszniewska-Matyszkiel [2011:  Thm.1] is met, and we see $V$ defined in (11) is the value function of the Bellman equation in DP, and $\tau^{*}$ maximizes $J$ in (2).
 ∎
-7 Conclusion
-This paper presents an alternative solution to avoid the prediction of the previous work: if the incomes of infinitely lived individuals are unobservable, efficient allocations are achieved only at the cost of invoking permanent inequality, leading to an immiserization of society.
-The proposed inter-period transfer mechanism achieves within-period full insurance by postponing risks.
-It does not trade off equal opportunities to become wealthy.
-It could be sustained by intergenerational cooporation.
-The result sheds light on efficient resource allocation for sustainable societies with equal opportunities.

### v16

@@
-Insuring uninsurable income
+
 Abstract
 We study dynamic mechanism design in a pure-exchange economy with privately observed idiosyncratic income.
-Classic hidden-income contracts attain constrained efficiency only at the cost of immiseration (Green 1987; Thomas–Worrall 1990).
+In the standard infinitely lived hidden-income benchmark of Green (1987) and Thomas–Worrall (1990), constrained-efficient allocations exhibit immiseration.
 We propose a simple recursive mechanism—adapted from Marcet–Marimon (1992)—that shifts each income shock forward by one period, keeps promised utilities in a bounded set, and, under a transparent “moderate risk-aversion” condition, delivers sequential efficiency.
-In a stationary overlapping-generations setting, we further provide an explicit condition on the initial promise that ensures budget sustainability; early cohorts pre-fund intertemporal smoothing so that every cohort attains a higher expected lifetime utility than under autarky.
+In a stationary overlapping-generations setting, we further show that under additional symmetry and curvature assumptions, a second-order approximation yields a sufficient condition for period-by-period budget balance; early cohorts pre-fund later transfers; for suitable initial promises, all cohorts are better off than under autarky.
 Our analysis uses a single state (promised utility), closed-form transfers, and a Bellman verification.
@@
 [^2] Earlier papers sometimes use immiserization[e.g., Phelan, 1998, Zhang, 2009].
-Recent work with persistent private information, Bloedel et al.
-[2025], also sharpens the mechanism of immiseration and explains the balance of the cost of incentive provision and backloading of high-powered incentives.
-When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but the knife-edge offers little comfort for moderate discounting.
+When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but recent work with persistent private information, Bloedel et al.
+[2025], shows that immiseration remains robust beyond the i.i.d.
+benchmark.
 A different route was proposed by Marcet and Marimon [1992].
@@
 Yet the mechanism has not been explored in a pure-exchange economy without investment or storage, where the scope for avoiding downward drift is a priori unclear.
-This paper adapts the Marcet–Marimon risk-shifting mechanism to an exchange economy in which agents cannot save.
+This paper adapts the Marcet–Marimon risk-shifting mechanism, that is, the current private shock changes tomorrow’s promised utility rather than today’s consumption, to an exchange economy in which agents cannot save.
 Promised utility is the sole state variable, so the model yields a closed-form characterization.
-If individual assets were allowed, one would require the set-valued approximation methods of Sleet and Yeltekin [2016]—a contrast confirmed by our closed-form example in Section 4.
-The contribution of my paper is threefold.
-(i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohort pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts.
+Relative to Green [1987] and Thomas and Worrall [1990], this paper changes three ingredients: agents have stochastic finite lifetimes in a stationary OLG environment, the mechanism is restricted to full within-period insurance, and incentives are provided through continuation value alone.
+The paper studies the consequences of that restricted class of mechanisms.
+The contribution of this paper is threefold.
+(i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohorts pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts for suitable initial promises.
+Unlike Fernandes and Phelan [2000], who enlarge the recursive state because serial correlation in private shocks destroys the one-dimensional promised-utility representation, this paper retains an i.i.d.
+hidden-income environment but identifies a mechanism in which current private shocks are absorbed by continuation promises.
+Under Assumption 1, this yields a simple one-state Bellman representation with full within-period insurance and bounded promised utility.
+Farhi and Werning [2007] and Phelan [2006] reinterpret and extend the classic hidden-information benchmark by replacing an infinitely lived individual with a dynasty of altruistically linked one-period-lived agents.
+In that dynastic literature, the central issue is how welfare, or promised utility, is inherited across parent and child under the incentive logic of Green (1987) and Thomas–Worrall (1990).
+By contrast, this paper studies a stationary overlapping-generations environment with stochastic survival.
+Its mechanism provides incentives by shifting the current private shock into next period’s promised utility; under Assumption 1, this is consistent with full within-period insurance.
+Demange [2002] studies Pareto properties of equilibrium risk sharing in stochastic OLG economies and identifies conditions under which adding a PAYG system at equilibrium is not Pareto-improving.
+By contrast, this paper studies a planner’s sequentially incentive-compatible transfer mechanism under hidden income.
+The mechanism is not a pure equilibrium PAYG reform: it uses continuation promises and, in the OLG extension, can involve transition pre-funding by early cohorts.
+Hence, the two papers address different institutional questions and need not deliver the same welfare conclusion.
 The organization of the paper is as follows.
@@
 The single state variable is the promised utility $\overline{v}_{1}(\lambda)$, which is strictly increasing in $\lambda$.
+We use $\lambda_{t}$ as a dual representation of promised utility $w_{t}=\overline{v}_{1}(\lambda_{t})$; all results could equivalently be written in $w_{t}$.
 2.2 Symmetric-information benchmark
@@
 The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
+Boundedness of $u$ is used later in the proof of Theorem 2 to show that the promised-utility martingale is bounded and hence convergent.
 Without specifying $\overline{C}$, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines $\overline{C}$ as well.
@@
 Given an initial promise $\lambda_{t}$ and any admissible transfer rule $\tau\in\mathcal{T}$, each individual’s discounted lifetime utility is
-$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
+$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{k}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
 This function will be used in Section 5 to verify that our mechanism solves the associated Bellman equation.
@@
 In general, a mechanism $\Gamma$ that arranges transfers $\{\tau_{t}\}_{t=0}^{\infty}$ needs to satisfy sequential incentive compatibility defined below.
-Definition 1.
-[Sequential incentive compatibility, Marcet and Marimon, 1992] Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
+Definition 1 (Sequential incentive compatibility).
+Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
 $(1-\beta)u(e+\tau(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\geq(1-\beta)u(e+\tau(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\quad\forall e,\tilde{e}\in E.$
 We further evaluate a mechanism according to the following notion.
-Definition 2.
-[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be a sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequentially incentive compatible mechanism.
+Definition 2 (Sequentially efficient mechanism).
+Fix a current state $(\lambda,e)$.
+Let $\mathcal{C}(\lambda,e)$ be the set of feasible sequentially incentive-compatible continuation mechanisms from that state.
+For any $\Gamma\in\mathcal{C}(\lambda,e)$, let
+$V_{1}^{\Gamma}(\lambda,e)$
+denote the individual’s expected discounted continuation utility, and let
+$V_{2}^{\Gamma}(\lambda,e)$
+denote the planner’s expected discounted continuation payoff (equivalently, minus the expected discounted transfer cost). We say that $\Gamma^{\prime}$ Pareto dominates $\Gamma$ at $(\lambda,e)$ if
+$V_{1}^{\Gamma^{\prime}}(\lambda,e)\geq V_{1}^{\Gamma}(\lambda,e),\quad V_{2}^{\Gamma^{\prime}}(\lambda,e)\geq V_{2}^{\Gamma}(\lambda,e),$
+with at least one strict inequality.
+A mechanism $\Gamma$ is sequentially efficient at $(\lambda,e)$ if it is sequentially incentive compatible and is not Pareto dominated at $(\lambda,e)$ by any other sequentially incentive-compatible continuation mechanism.
+Remark 2.
+In this paper, Pareto comparisons are made over the continuation payoff pair $(V_{1},V_{2})$, conditional on the current state $(\lambda,e)$.
+Accordingly, for any fixed state $(\lambda,e)$, any solution to problem (1) yields a Pareto-efficient continuation payoff pair in this sense.
 In addition to taking these things into account, since it is known to lead to an immiseration outcome, the planner does not hope to manipulate the transfer function $\tau^{*}$.
@@
 $\geq(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\text{ for }e,\tilde{e}\in E.$ (4)
-Such a promise $\lambda^{\prime}$ that satisfies both (3) and (4) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$.
-By a feasible $\lambda^{\prime}$ we mean that the updated promise keeps the level of expected utility for the next period below the supremum of each individual’s utility function.
-That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$.
-For given $\lambda>0$, if $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))<S$, define $\lambda^{\prime}$ such that
-$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))\big{)}.$ (5)
-Under the updating rule of $\lambda^{\prime}$ in (5), misreporting never raises the weighted sum $\lambda u(e+\tau^{*}(e|\lambda))+v_{2}(\lambda,e)$, hence truth-telling is optimal.
+To construct a continuation promise satisfying both (3) and (4), define the deviation of the planner’s continuation payoff from its average by
+$\Delta v_{2}(\lambda,e):=v_{2}(\lambda,e)-\overline{v}_{2}(\lambda).$
+Since $v_{2}$ is the discounted value of $\tau$, a larger value of $\Delta v_{2}(\lambda,e)$ corresponds to a smaller current transfer (or a larger current contribution).
+We choose $\lambda^{\prime}(e\mid\lambda)$ so that this deviation is exactly offset by a change in promised continuation utility:
+$\beta\lambda\bigl(v_{1}(\lambda^{\prime}(e\mid\lambda))-v_{1}(\lambda)\bigr)=\Delta v_{2}(\lambda,e).$
+Equivalently,
+$\lambda\Bigl[(1-\beta)u(e+\tau^{*}(e\mid\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e\mid\lambda))\Bigr]$
+$=\lambda\Bigl[(1-\beta)u(e+\tau^{*}(e\mid\lambda))+\beta\overline{v}_{1}(\lambda)\Bigr]+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda).$
+Hence, the continuation payoff under truthful reporting coincides with the weighted-sum problem used in (1). Moreover, since
+$E[\Delta v_{2}(\lambda,e)\mid\lambda]=0,$
+taking expectations yields the promise-keeping condition.
+The remaining issue is feasibility. Since $\overline{v}_{1}$ is bounded above by
+$S:=\sup_{x\in\mathbb{R}_{+}}u(x),$
+the inverse $\overline{v}_{1}^{-1}$ is well defined only if the target continuation utility remains below $S$. If
+$\overline{v}_{1}(\lambda)+\frac{1}{\beta\lambda}\bigl(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda)\bigr)<S,$
+where $e^{M}$ is the maximal income realization, then $\lambda^{\prime}(e\mid\lambda)$ is well defined for every $e\in E$, and we may define
+$\lambda^{\prime}(e\mid\lambda)=\overline{v}_{1}^{-1}\!\left(\overline{v}_{1}(\lambda)+\frac{1}{\beta\lambda}\bigl(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)\bigr)\right).$ (5)
 On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (5) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the planner must find the closest promise for which there is a certainty equivalent to the promised level of lifetime utility.
@@
 Since the definition of $\lambda^{\prime}$ in (6) is an extension of (5), we can use it for the case where $\lambda^{\prime}$ could also be defined in (5). We define a mechanism like this.
-Definition 3.
-(Risk-shifting mechanism) A mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6).
-Within each period $t$, consumption is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1})$; schocks affect only $\lambda_{t+1}$.
 A sufficient condition for the sequence of promise $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (5) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal consumption $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$.
@@
 Interpretation.
-Arrow-Pratt risk aversion at the contract consumption level is smaller than a scaled income spread; the condition is mild when $\beta$ is close to one.
-Under Assumption 1, feasibility does not bind in the promise-update map, and the recursive rule (5) applies period by period.
+Assumption 1 compares Arrow-Pratt absolute risk aversion at the contract consumption level with a scaled income spread.
+A convenient class of examples is given by utility functions with nonincreasing absolute risk aversion, since in that case the condition is easiest to verify at low consumption levels.
+In Section 4 our numerical specification
+$u(c)=\frac{(c+1)^{1+\gamma}}{1+\gamma},\qquad\gamma=-3,$
+has
+$\frac{u^{\prime\prime}(c)}{u^{\prime}(c)}=\frac{3}{c+1},$
+so absolute risk aversion is decreasing.
+We define risk-shifting mechanism as follows.
+Definition 3.
+(Risk-shifting mechanism) A risk-shifting mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6).
 Remark 3.
-We could remove condition Assumption 1 by removing the assumption of bounded utility; dropping it would not violate sequential efficiency but would complicate the martingale proof.
-Boundedness is used to ensure that $\{\overline{v}_{1}(\lambda_{t})\}$ converges.
+Under Assumption 1, the feasibility correction never binds, so current consumption is constant at
+$c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1}),$
+and the realized shock affects only the next-period promise.
+If the feasibility correction (6) binds, then agents whose unconstrained continuation promise would exceed the maximal feasible utility level receive the capped continuation promise instead.
+Since their contribution is capped at the threshold level, they retain part of the current income shock in current consumption.
+Hence, full within-period insurance fails for those realizations, but the promise-keeping condition holds weakly in discounted lifetime utility.
+Remark 4.
+If utility were unbounded above, the feasibility issue behind Assumption 1 would disappear, but the martingale-convergence argument would then require a different integrability condition.
 Lemma 1.
@@
 3.1 Individual level: bounded promised utilities
-This subsection shows that the sequence of promised utilities forms a bounded martingale and therefore converges, so immiseration cannot occur.
+This subsection shows that the sequence of promised utilities forms a martingale and therefore converges, so immiseration cannot occur.
 Let $(\mathcal{F}_{t})$ be the natural filtration generated by $\{\lambda_{s}\}_{s\leq t}$.
@@
 Theorem 2.
-The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a bounded martingale and converges to an integrable random variable a.s.
+The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a martingale and converges to an integrable random variable a.s.
 (no immiseration).
 Proof: see Appendix A.
-Because promised utility is a martingale with bounded support, the mechanism is a fair intertemporal lottery: expected lifetime utility is preserved ex-ante, but its realisation drifts neither up nor down on average.
+We have for $n\in\mathbb{N}$ $\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e.
+$[P]$.
+Hence, $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale, and promised lifetime utility is preserved in conditional expectation over time.
+Realized promised utility may vary across histories, but it exhibits no systematic drift.
+Since the process is bounded, it converges almost surely to an integrable random variable, so immiseration does not occur.
+In Thomas and Worrall [1990], the state variable (promised utility) $V^{t}$ converges to -$\infty$, whereas the martingale object is the lender’s marginal value process $U^{\prime}_{*}(V^{t})$.
 3.2 Planner’s budget and sustainability
@@
 As we see in the following proposition, under a second-order (Taylor) approximation, if incomes have a symmetric distribution, a sufficient condition for the budget balance in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
-Theorem 3.
+Proposition 1 (second-order sufficient condition for budget balance).
 In a stationary overlapping-generations society with i.i.d. income risk, if
@@
 then planner’s period-by-period balance $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ under a second-order (Taylor) approximation.
-Proof: Second-order approximation shows decreasing differences. See Appendix C.
+Proof sketch.
+The key step is to show that the second-order approximation of the average-transfer increment $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is decreasing in $n$.
+For a fixed current promise $\lambda_{n-1}$ and realization $e$, define
+$\Delta_{n-1}:=\overline{v}_{1}(\lambda_{n}(e\mid\lambda_{n-1}))-\overline{v}_{1}(\lambda_{n-1})=\frac{1-\beta}{\beta\lambda_{n-1}}(e-E[e_{n}]).$
+We expand
+$u^{-1}(\overline{v}_{1}(\lambda_{n}(e\mid\lambda_{n-1})))-u^{-1}(\overline{v}_{1}(\lambda_{n-1}))$
+around $\overline{v}_{1}(\lambda_{n-1})$.
+As shown in Appendix C, a third-order Taylor expansion yields a second-order approximation for $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ after taking expectations.
+In the symmetric case, the odd-order term cancels in expectation, so the leading term is the second-order term.
+Assumptions (ii) and (iii), together with Jensen’s inequality, imply that this second-order term is decreasing in $n$.
+Hence the second-order approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\geq 1}$ is decreasing.
+Therefore,
+$\overline{\tau}_{n}-\overline{\tau}_{n-1}\leq\overline{\tau}_{1}-\overline{\tau}_{0}\qquad\text{(under the second-order approximation),}$
+and substituting this bound into
+$g_{t}(\lambda_{0})=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
+shows that condition (iv) is sufficient for $g_{t}(\lambda_{0})\geq 0$ for all $t\in\mathbb{N}$. See Appendix C for the complete proof.
 4 Numerical examples
-An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Theorem 3 is given by the second-order polynomial approximation as follows.
+An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Proposition 1 is given by the second-order polynomial approximation as follows.
 Corollary 1.
-The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Theorem 3 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
+The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Proposition 1 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
 $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))\geq\sqrt{\frac{(1-\alpha r)^{2}}{\alpha r^{2}(1-\alpha)}\mathrm{Var}[e_{t}]},$ (9)
@@
 If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (9) may be unnecessarily strict.
-We illustrate the mechanism with a CRRA utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
+We illustrate the mechanism with a Laurent polynomial utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
   • Utility $u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma}$, $\gamma=-3$
@@
   • Initial state $\lambda_{0}=50\in[\lambda_{CE},133]$
-The values follow Corollary 1: the lower bound $\lambda_{CE}=1/u^{\prime}\bigl{(}E[u(e_{t})]\bigr{)}=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
+The values follow Corollary 1: in this calibration the lower bound $\lambda_{CE}=1/u^{\prime}\bigl(u^{-1}(E[u(e_{t})])\bigr)=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
 Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals of the same generation and corresponding transfer balances for $100$ periods for the above case.
@@
 The random numbers are generated with the same symmetric distribution $P_{s}$.
-The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and supports Theorem 3.
+The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and is consistent with Proposition 1.
 Figure 3: Sample means of budget balance $\{g_{t}(\lambda_{0})\}$
-Remark 3.
-Classic hidden-income contracts [e.g.
-Green, 1987, Thomas and Worrall, 1990] spread shocks within each period: truth telling is rewarded by a higher spot transfer, so lifetime utilities diverge over time and immiseration follows.
-Under our mechanism, by contrast, the realized income $e_{t}$ only determines next period’s promise $\lambda_{t+1}$; consumption in period $t$ is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda^{-1}_{t})$ with probability $1$.
-The shock is therefore shifted one step forward—a transfer of risk to the future.
-This intertemporal insurance does make the average transfer profile $\{-\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ fall with age, threatening the planner’s long-run budget.
-Theorem 3 shows the problem can be neutralized by choosing an initial promise $\lambda_{0}$ that gives early cohorts negative net transfers, so later positive transfers are prefunded.
-Finally, note a key difference from Marcet and Marimon [1992]: in their growth model, investment returns are themselves the incentive, and the contract is bilateral and non-competitive, which can depress the manager’s utility below autarky.
-Our pure-exchange setting has no such external scale effect, so every cohort attains a higher lifetime utility than autarky.
 5 Dynamic-programming verification
 This section verifies that the transfer rule $\tau^{\ast}$ solves the infinite-horizon Bellman equation and that $V$ is the corresponding value function (under condition (8) of Lemma 1).
+For a given current promise $\lambda$, let $\Lambda(\lambda)$ denote the set of measurable controls $\tau:E\to\mathbb{R}$ such that the implied continuation promise $\lambda^{\prime}_{\tau}(e\mid\lambda)$ defined by (5) whenever feasible, and by (6) otherwise, is well defined for all $e\in E$, and the pair $(\tau,\lambda^{\prime}_{\tau})$ satisfies the promise-keeping condition (3) and the sequential incentive constraint (4).
 The Bellman equation (DP) is given by
-$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)]\},$ (10)
-where $\Lambda(\lambda_{t})$ is a set of measurable controls satisfying incentive and promise-keeping, and $\lambda_{t+1}$ is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{R}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (6).
+$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda^{\prime}_{\tau}(e_{t}|\lambda_{t}),t+1)]\}.$ (10)
+Under Assumption 1, $\lambda^{\prime}_{\tau}$ is given by (5).
 Our candidate solution is $V$ in (11) with control $\tau^{\ast}$.
+[^4] Although we write $V(\lambda_{t},t)$ to match the terminal-condition formulation of Wiszniewska-Matyszkiel [2011], the economic environment is stationary; the time argument is used only for the dynamic-programming verification.
+Along the candidate mechanism, $V(\lambda_{t},t)=v_{1}(\lambda_{t})$.
 $V(\lambda_{t},t)=\mathbb{E}_{t}\big[u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big],\quad\text{for }\,\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}.$ (11)
@@
 6 Conclusion
-This paper shows that the grim ”immiseration“ prediction of Green [1987] and Thomas and Worrall [1990] is not inevitable.
-A simple risk shifting rule that postpones risk can deliver near full within-period insurance while keeping promised utilities bounded.
-Provided the initial promise satisfies a transparent inequality, inter-generational cooperation sustains the mechanism, and every cohort attains higher lifetime utility than under autarky, with no within-cohort inequality.
-Appendix A Proof of propositions 1 and 2
+This paper shows that immiseration is not unavoidable within the risk-shifting mechanism studied here.
+The mechanism induces a bounded promised-utility process and, under the paper’s sufficient condition, is sequentially efficient.
+In the stationary OLG extension, the paper also identifies sufficient conditions under which intergenerational transfers sustain the mechanism and can improve on autarky.
+Appendix A Proof of theorems 1 and 2
 Proof of Lemma 1.
@@
 Proof of Theorem 1.
-With respect to $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ defined throughout in (5), the propositional statement is shown by Marcet and Marimon [1992]. For completeness, we report the entire proof.
+Similar proof can be found in Marcet and Marimon [1992].
 We first show that $\mathcal{M}$ is sequentially incentive compatible. For $\lambda_{t}>0$ with $F(\lambda_{t})<S$
@@
 The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) (planner’s static problem at given $\lambda$) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
-We proceed to show that if condition (8) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms.
+We next show that, under condition (8), the continuation rule induced by $\mathcal{M}$ solves problem (1) at every feasible current state.
 By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible.
-By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (5).
+By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined by (5).
 Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period.
-Therefore $\mathcal{M}$ is a Pareto optimal transfer mechanism.
+Therefore, conditional on the current promised state and within the admissible class where (5) applies, $\tau^{*}$ solves the planner’s static problem at each date and induces a Pareto-efficient continuation payoff pair.
 It remains to prove that under condition (8) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism.
-Suppose, contrary to our claim, that there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$.
-Let $(v_{1}^{*},v_{2}^{*})$ be the present value achieved by $\Gamma$.
-Set $\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*})$.
-Since condition (8) holds for any $\lambda>0$, this $\lambda_{0}$ satisfies $F(\lambda_{0})<S$ as in the proof of Lemma 1.
-We may now use $(\lambda_{0},e)$ as the initial condition for $\mathcal{M}$ under condition (8).
-Then, by construction, each individual has the same present value for both contracts.
-Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$.
-However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under the condition (8).
-∎
+Suppose, contrary to our claim, that under Assumption 1 there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$.
+Let $(v_{1}^{*},v_{2}^{*})$ denote the continuation payoff pair induced by $\Gamma$.
+Since $v_{1}^{*}$ lies in the range of $\overline{v}_{1}$, define
+$\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*}).$
+Because condition (8) depends only on the primitives and holds for every $\lambda>0$, it also holds at $\lambda_{0}$.
+Hence $\mathcal{M}$ is well defined from the initial state $(\lambda_{0},e)$, and its continuation rule is given by (5).
+By construction, $\Gamma$ and $\mathcal{M}$ deliver the same continuation utility to the agent, so $v_{1}^{*}=\overline{v}_{1}(\lambda_{0})$.
+Since $\Gamma$ Pareto dominates $\mathcal{M}$, it must satisfy
+$v_{2}^{*}>v_{2}(\lambda_{0},e).$
+Hence
+$\lambda_{0}v_{1}^{*}+v_{2}^{*}>\lambda_{0}\overline{v}_{1}(\lambda_{0})+v_{2}(\lambda_{0},e),$
+which contradicts the optimality of $\tau^{*}$ in problem (1) at $(\lambda_{0},e)$. ∎
 Proof of Theorem 2.
@@
 $=\overline{\tau}_{n}\quad\text{ for }n=2,3,\cdots.$
-Appendix C Proof of Theorem 3
-The proof of Theorem 3 uses the following lemma.
+Appendix C Proof of Proposition 1
+The proof of Proposition 1 uses the following lemma.
 Lemma C.1.
-Under the assumptions (i), (ii), (iii) in Theorem 3, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
+Under the assumptions (i), (ii), (iii) in Proposition 1, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
 Proof.
 First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (5).
+We use a third-order Taylor expansion to obtain the second-order approximation used in the proof of Proposition 1.
 Using the third-order Taylor polynomial, for a given promise $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
@@
 $=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big(u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
-It follows that the third order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence.
+It follows that the second-order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ obtained from the third-order expansion is a decreasing sequence.
 We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (6). We change $\Delta_{n-1}$ in (C.1) to
@@
 By a similar argument, the second-order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
-We prove Theorem 3 below.
-Proof of Theorem 3.
+We prove Proposition 1 below.
+Proof of Proposition 1.
 From Lemma C.1, with the second-order polynomial approximation, we have
@@
 $\lambda_{t}\big[(1-\beta)$ $u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big]$
-$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\overline{v}_{2}(\lambda_{t})$
-and we see that the last terms are maximized by $\tau^{*}$. It follows that (E.1) is written as
-$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)],$
-and we see the desired result.
+$=\lambda_{t}(1-\beta)u(e_{t}+\tau(e_{t}\mid\lambda_{t}))+v_{2}(\lambda_{t},e_{t})+\lambda_{t}\beta\overline{v}_{1}(\lambda_{t})-\overline{v}_{2}(\lambda_{t}).$
+Thus, up to the additive constant $\lambda_{t}\beta\overline{v}_{1}(\lambda_{t})-\overline{v}_{2}(\lambda_{t})$, the Bellman integrand is exactly the objective of problem (1).
+Hence $\tau^{*}$ attains the supremum.
+Therefore $\tau^{*}$ maximizes (E.2).
 The task is now to check the terminal condition.